### v1

# Convex Relaxations of ReLU Neural Networks Approximate Global Optima in Polynomial Time

Abstract
In this paper, we study the optimality gap between two-layer ReLU networks regularized with weight decay and their convex relaxations. We show that when the training data is random, the relative optimality gap between the original problem and its relaxation can be bounded by a factor of $O(\sqrt{\log n})$, where $n$ is the number of training samples. A simple application leads to a tractable polynomial-time algorithm that is guaranteed to solve the original non-convex problem up to a logarithmic factor. Moreover, under mild assumptions, we show that with random initialization on the parameters local gradient methods almost surely converge to a point that has low training loss. Our result is an exponential improvement compared to existing results and sheds new light on understanding why local gradient methods work well.
1 Introduction
After the tremendous success of deep learning [1], data-driven approaches have become a prominent trend in various areas of computer science. Perhaps a surprising fact is that despite the highly non-convex landscape of deep learning models [2], local gradient methods such as stochastic gradient descent (SGD) [3] or ADAM [4] find nearly-global minimizers of the network extremely well. This mystery has gained wide attention in the learning theory community, and many have worked on the problem of proving convergence results of local methods under certain assumptions such as the infinite-width limit [5], heavy overparametrization [6], [7],[8], [9], and milder width assumptions [10].
Figure 1: Convex relaxations of different widths. Here, we show the optimal value of the relaxed problem for different numbers of subsampled hyperplane arrangement patterns and regularization. Note that $n=300,d=10$, hence there are $\approx 30^{10}$ variables in the convex reformulation. However, using $\approx 30$ neurons in the relaxed problem optimizes the objective well.
In contrast to the empirical success of local gradient methods, training a simple two-layer neural network with ReLU activation is proven to be NP-Hard [11]. The sharp contrast between existing works on trainability guarantees and impossibility results occurs as an increase in the width $m$, though it may seem to make the problem harder, actually makes it easier. Hence, it is a natural interest to understand the critical width $m^{*}$ that guarantees polynomial time algorithms where $m\geq m^{*}$, and the corresponding algorithm.
Recently, it was shown that two-layer and deeper ReLU networks have exact convex reformulations [12, 13]. However, the number of variables in these convex problems can be exponentially large. For a critical width $m^{*}\leq n+1$, when $m\geq m^{*}$, we have an algorithm that exactly solves the non-convex problem with complexity polynomial in $n$ (the number of data) and exponential in $r$ (the rank of the dataset). Such an algorithm is given by considering the convex reformulation of the original problem which has approximately $(n/r)^{r}$ variables and solving the reformulated problem with standard interior-point solvers. Due to the exponential factor in $r$, exactly solving the reformulated problem is intractable, and [12] proposed to solve the randomly relaxed version of the exact reformulation with $m\ll(n/r)^{r}$ variables, where $m$ is only polynomially large (see Section 2.1 for details). Though the relaxed problem has polynomial complexity with respect to all dimensions, and works well in practice (see Figure 1), the optimality gap between the relaxed and the original problem has remained unknown. In Figure 1, width corresponds to the number of variables that the convex relaxation has.
In this paper, we study the randomized relaxation of the exact convex reformulation and find that:
  • Under certain assumptions on the input data, the relative optimality gap between the non-convex problem and the randomized relaxation is bounded by a logarithmic factor. This shows that the convex relaxations run in polynomial-time and have strong approximation properties.
  • Using results from [14] and [15], we show that with high probability, local gradient methods with random initialization converges to a stationary point that has $O(\sqrt{\log n})$ relative training error with respect to the global minimum.
  • We propose a tractable polynomial-time algorithm which is polynomial in all dimensions that can approximate the global optimum within a logarithmic factor. We are not aware of any similar result for regularized ReLU networks.
The paper is organized as follows: in Section 1.1, we go through a brief overview of related works. Section 2 is devoted for a detailed description of the contributions we made. Section 3 and Section 4 illustrates the overall proof strategy, novel approximation guarantees, and the geometry behind it: we first obtain approximation guarantees of an “easier version” of the randomly relaxed problem, and then solve the relaxed problem using the result. We wrap the paper with Section 5, presenting possible further discussions.
1.1 Prior and Related Work
Convex reformulation of neural networks: Starting from [12], an extensive line of work have discussed convex reformulation of neural networks. The main idea that lies in this line of work is that for different neural network architecture, e.g. for CNNs [16], transformers [17], multi-layer networks [18], vector outputs [19], etc, we have an exact convex reformulation with a different regularizer. Hence, neural network architectures can be understood as imposing different regularization to the convex problem.
Moreover, these convex reformulation have shown to give interesting insights on the training of neural networks. For example, [14] characterizes all stationary points of a two-layer neural network via global optimum of corresponding convex problems, and [20] discusses the intrinsic complexity of training two-layer networks with the view of convex duality and equivalence with the MAX-CUT problem.
Training complexity of two-layer neural networks: We use $n$ to denote the number of data, $d$ the dimension of data, and $m$ to be the width of the model. Several papers have discussed the theoretical complexity of training simple ReLU networks, and exact methods to train them. It is proven that for single-width neural networks, it is NP-Hard to train a two-layer ReLU network [11], and even approximating the optimal error with $(nd)^{1/poly(\log\log(nd))}$ is proven to be NP-Hard [21]. [22] gives an approximate algorithm to solve the single-width neural network problem with $O(n^{k})$ complexity and $n/k$ approximate error for general input, and constant error bound in the student-teacher setting.
Some exact algorithms to train two-layer network have been discussed, though due to the inherent complexity of training two-layer networks, these algorithms are often intractible in practice. [23] presents an algorithm with $O(2^{m}n^{dm}poly(n,d,m))$ complexity, where $m$ is the width of the network. [24] shows that when both the input and weights are constrained on a unit ball, we can train a two-layer network to have $\epsilon$ error bound with $O((2^{m/\epsilon})^{O(1)}n^{O(1)})$ complexity. Training guarantees in the student teacher setting have also been widely discussed: [25] discusses gaussian inputs and in the student teacher setting, we can exactly obtain the teacher parameters in $poly(n)$ time, and [26] discusses the same problem and propose an algorithm that works in $poly(n,m,d)$.
Note that the aforementioned complexity results are on training problems without regularization. When we have an $l_{2}$ regularized problem, the best exact solution to the problem is presented in [12], where they propose an algorithm of $O(d^{3}(n/d)^{3d})$ complexity provided that the width of the network is sufficiently large. Extending the idea, [27] proposes an approximate method with complexity $O(d^{2}m^{2})$ that works for $m\geq n/\xi$ for a predetermined error threshold $\xi$. However, their results require a very wide neural network, and the relative optimal error bound is unclear.
Trainability in the mild overparametrization regime: In the mild overparametrization regime where the number of parameters are $\tilde{O}(n)$, training guarantees are not well estabilished. Some noteworthy works are [28] that shows when the number of parameters for a single layer exceeds $n$, differentiable local minima are global, and [10], which shows that for certain datasets only a polylogarithmic width is enough to provide training guarantees.
For useful lemmas that will be used throughout the paper, see Appendix A.
2 Main Results
2.1 Preliminaries
First, we go through a formal description of the optimization problems that we are interested in. Let the data matrix $X\in\mathbb{R}^{n\times d}$, the label vector $y\in\mathbb{R}^{n}$, and weight decay regularization $\beta>0$. Consider the training problem
$p^{*}:=\min_{u_{j},\alpha_{j}}\frac{1}{2}\lVert\sum_{j=1}^{m}(Xu_{j})_{+}\alpha_{j}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{j=1}^{m}(\lVert u_{j}\rVert_{2}^{2}+\lVert\alpha_{j}\rVert_{2}^{2}).$ (1)
The convex counterpart of problem (1) can be written as in [12]:
$p_{0}^{*}:=\min_{u_{i},v_{i}\in\mathcal{K}_{D_{i}}}\frac{1}{2}\lVert\sum_{D_{i}\in\mathcal{D}}D_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i\in\mathcal{I}}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}).$ (2)
Convex reformulation: Let’s say $[n]=\{1,2,...,n\}$ and $D_{i}$ from $i\in[P]$ denote all possible hyperplane arrangement patterns $diag(\mathbbm{1}[(Xu)_{+}])$, where $u\in\mathbb{R}^{d}$. Also, $\mathcal{K}_{D}=\{u\ |\ (2D-I)Xu\geq 0\}$ are cones where each optimization variable are constrained at. When $\mathcal{D}$ is chosen to be $\mathcal{D}=\{D_{i}\ |\ i\in[P]\},\{u_{i}^{*},v_{i}^{*}\}_{i=1}^{P}$ are global minima of problem (2),
$m^{*}=\sum_{i=1}^{P}\mathbbm{1}[u_{i}^{*}\neq 0]+\mathbbm{1}[v_{i}^{*}\neq 0],$
[…]
$\mathcal{D}=\{\tilde{D}_{i}=diag(\mathbbm{1}[Xg_{i}\geq 0])\ |\ g_{i}\sim\mathcal{N}(0,I_{d}),i\in[\tilde{P}]\},$
i.e. instead of using “every possible” hyperplane arrangement pattern, we randomly sample $\tilde{P}$ patterns given by a random Gaussian vector. Note that we are using a very small portion of the set of all hyperplane arrangement patterns, and it is not trivial that we will get good approximation guarantees.
Next, we illustrate an assumption on the data distribution for the main results.
Assumption on Dataset: Throughout the paper, we assume that the data distribution follows $X_{ij}\sim\mathcal{N}(0,1)$ i.i.d., $n/d=c$, where $c$ is a constant, and denote as (A1). (A1) can be extended to data that are isotropic and subgaussian, hence a bounded isotropic dataset can have similar proof strategies and theoretical results as this paper. Here, we assume Gaussianity for simplicity. Also, the regime $n/d=c$, i.e. $n\asymp d$, is extensively studied in the literature [30], [31], [32]. The given structure makes quantitive comparison between the randomized relaxation and the original problem with the lens of random matrix theory, and provides better bounds than arbitrary inputs.
2.2 Overview of Theoretical Results
We are now ready to state the main results of the paper. The first result is the optimality bound between the non-convex problem (1) and its Gaussian relaxation, assuming (A1).
Theorem 1.
(Informal) Consider the two-layer ReLU network training problem
$p^{*}:=\min_{u_{j},\alpha_{j}}\frac{1}{2}\lVert\sum_{j=1}^{m}(Xu_{j})_{+}\alpha_{j}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{j=1}^{m}(\lVert u_{j}\rVert_{2}^{2}+\lVert\alpha_{j}\rVert_{2}^{2}),$
and its convex relaxation
$\tilde{p}^{*}:=\min_{u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}}\frac{1}{2}\lVert\sum_{\tilde{D}_{i}\in\mathcal{D}}D_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i\in\mathcal{I}}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}).$ (3)
Here, $|\mathcal{D}|=m/2$ and the elements are sampled by the hyperplane arrangement patterns of random Gaussian vectors. Assume (A1), $d$ is sufficiently large, and suppose that $m$ satisfies $m\geq\max\{m^{*},320(\sqrt{c}+1)^{2}\log(\frac{n}{\delta})\}$. Then, with high probability,
$p^{*}\leq\tilde{p}^{*}\leq C\sqrt{\log(2n)}\ p^{*}.$
for some constant $C\geq 1$.
Remark.
To the best of our knowledge, the above result provides the first polynomial-time approximation guarantee for regularized ReLU NNs. Also note that typically we have $p^{*}\rightarrow 0$ as $\beta\rightarrow 0$, implying $\tilde{p}^{*}-p^{*}\rightarrow 0$.
For a detailed statement of the theorem, see Appendix D. A few points on Theorem 1 is worth mentioning. First, it gives a guarantee of the convex relaxation of the original problem when we sample only $\max\{m^{*},O(\log(n))\}$ hyperplane arrangement patterns. This is an exponential improvement over existing convex reformulations, and as $m^{*}$ is much smaller than $n+1$ in practice, the width bound for Theorem 1 is practical. Also, with such characterization, we obtain a polynomial-time approximate algorithm that works with guarantees by simply solving the convex relaxation with standard interior-point solvers [33].
Theorem 2.
[…]
Moreover, with the characterization of stationary points in [14], we can prove that local gradient methods converge to “nice” stationary points under the assumption below:
(A2) While local gradient method iterates, only a portion $p<1/2$ of the initial hyperplane arrangement patterns change until it converges.
Figure 2: Verification of (A2) for gradient descent. The upper figure shows how many hyperplane arrangement patterns change for random data, and the lower figure shows how many of them change for MNIST.
(A2) is also considered in [15], where they show that SGD with random initialization preserves most of the hyperplane arrangement patterns. The specific lemma, which they refer to as the coupling lemma, is as follows. It is also verified for gradient descent, both for random data and MNIST in Figure 2.
Lemma 1.
(Coupling lemma, [15]) With high probability over the random initialization that follows $\mathcal{N}(0,m^{-1}I)$, and suppose the data is generated from $l$ underlying distributions. For every $\tau>0$, $t=\tilde{O}(\tau/\eta)$, at least $1-e\tau l/\sigma$ portion of the hyperplane arrangement patterns remain the same.
The idea is that we can map each stationary point to a global minima of a convex problem, and as we initialized at random, we can think of the stationary points as global minimizers of the convex problem with randomized hyperplane arrangement patterns.
Theorem 3.
(Informal) Consider the training problem $\min_{u_{j},\alpha_{j}}\mathcal{L}(u,\alpha),$ where the loss function $\mathcal{L}$ is given as
$\mathcal{L}(u,\alpha)=\frac{1}{2}\lVert\sum_{j=1}^{m}(Xu_{j})_{+}\alpha_{j}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{j=1}^{m}(\lVert u_{j}\rVert_{2}^{2}+\alpha_{j}^{2}).$
Assume (A1),(A2) and $d,m$ are sufficiently large. For any random initialization $\{u_{i}^{0},\alpha_{i}^{0}\}_{i=1}^{m}$, suppose local gradient method converged to a stationary point $\{u_{i}^{\prime},\alpha_{i}^{\prime}\}_{i=1}^{m}$. Then, with high probability,
$\mathcal{L}(u^{\prime},\alpha^{\prime})\leq C\sqrt{\log(2n)}\mathcal{L}(u^{*},\alpha^{*}),$
for some $C\geq 1$. Here, $\{u_{i}^{*},\alpha_{i}^{*}\}_{i=1}^{m}$ is a global optimum of $\mathcal{L}(u,\alpha)$.
Remark.
The above result shows that there exists many stationary points whose objective value is a logarithmic approximation of the global optimum. Therefore, first-order optimizers such as SGD and Adam can approximate the global optimum even when they converge to stationary points.
Note that our analysis is not based on explicit iteration of local gradient methods, or does not exploit a NTK - based analysis. Rather, it follows from a simple yet clear analysis based on the lens of convex optimization.
2.3 Notations for the proof
Before discussing the proof strategy, we clarify some frequently used variables in this section. We use $D_{i}$ for $i\in[P]$ to denote all possible hyperplane arrangements, $P$ to denote the number of all possible hyperplane arrangements, $\tilde{D}_{i}$ to denote randomly selected hyperplane arrangement patterns, and $\tilde{P}$ the number of such subsamples. We also use $\mathcal{K}_{D}$ as in preliminaries. $\mathcal{M}$ is used to denote
$\mathbb{E}_{g\sim\mathcal{N}(0,I_{d})}[diag[\mathbbm{1}(Xg\geq 0)]XX^{T}diag[\mathbbm{1}(Xg\geq 0)]],$
and $\kappa$, which is analogous to the condition number, to denote
$\frac{\lambda_{max}(XX^{T})}{\lambda_{min}(\mathcal{M})},$
provided that $\mathcal{M}$ is invertible. At last, $m^{*}$ is used to denote the number of nonzero variables for the optimal solution of the convex reformulation.
2.4 Overall proof strategy
To prove Theorem 1, we first consider the unconstrained problem
$\tilde{p_{1}}^{*}:=\min_{u_{i},v_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{\tilde{D}_{i}\in\mathcal{D}}D_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i\in\mathcal{I}}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}).$ (4)
The convex problem (4) was first introduced in [34], where they first solve the unconstrained problem and decompose it into cone constraints to solve the original problem (3). The problem is further equivalent to
$\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{P}D_{i}Xw_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{P}\lVert w_{i}\rVert_{2},$ (5)
due to triangle inequality. We use a similar strategy that was introduced in [34]: first, we show that the unconstrained problem enjoys good approximation guarantees, even though we use only a subsample of the whole possible hyperplane arrangement patterns. After that, for global optimum $\{w_{i}^{*}\}_{i=1}^{m}$ of problem (5), we find $u_{i},v_{i}\in\mathcal{K}_{D_{i}}$ that satisfies $u_{i}-v_{i}=w_{i}^{*}$ and minimal norm sum $\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}$. We find that this decomposition does not increase the norm up to $O(\sqrt{\log(2n)})$ factor, finally leading to the guarantee to the original problem.
3 Guarantees for the Unconstrained Convex Relaxation
Let’s start with the unconstrained optimiization problem Equation 5. Note $\tilde{p}_{1}^{*}$ to be the optimal value of the random relaxation of the unconstrained problem, and $p_{1}^{*}$ to be the optimal value of the unconstrained problem using all possible hyperplane arrangement patterns. We wish to find constants $C_{1},C_{2}>0$ depending only with the data matrix $X$ that satisfies
$0\leq\tilde{p}_{1}^{*}-p_{1}^{*}\leq C_{1},\quad p_{1}^{*}\leq\tilde{p}_{1}^{*}\leq C_{2}p_{1}^{*}.$
As a warmup, we first give results when the problem is unregularized. We can show that $p_{1}^{*}=\tilde{p}_{1}^{*}=0$ with high probability when we sample sufficiently many hyperplane arrangement patterns. After that, we give approximation results for a gated ReLU problem with $l_{2}$ regularization and use them to prove that
$C_{1}=\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}},C_{2}=2\sqrt{2}\sqrt{\frac{\lambda_{\max}(XX^{T})}{\lambda_{\min}(\mathcal{M})}},$
holds with probability at least $1-\delta$, provided that we sample sufficiently many hyperplanes and the dimension $d$ is sufficiently large. We defer the proofs to Appendix B.
3.1 Warmup: Unconstrained Relaxation Without Regularization
To show that $p_{1}^{*}=0$, we prove that we can approximate any vector with $2n$ hyperplane arrangement patterns. The proof is consistent with general overparametrization arguments, which state that when the width of the network is $m\geq n$, any local minimum becomes the global minimum. Here, we carefully choose $2n$ hyperplane arrangement patterns where each arrangement pair $D_{1,i},D_{2,i}$ differ only at the $i$-th diagonal entry. With $n$ such pairs we can express any vector $y\in\mathbb{R}^{n}$ as a linear combination of vectors in the column space of $\{(D_{1,i}-D_{2,i})u|u\in\mathbb{R}^{n}\}=\{ke_{i}|k\in\mathbb{R}\}$, proving that there exists $u_{i}$ s satisfying
$\sum_{i=1}^{n}(D_{1,i}-D_{2,i})Xu_{i}=y,$
meaning that if we have all possible hyperplane arrangement patterns we can express any vector $y$ with zero error. The specific proposition is as follows.
Proposition 1.
Suppose no two rows of $X$ are parallel. There exists hyperplane patterns $\tilde{D}_{1}$, $\tilde{D}_{2}$, …, $\tilde{D}_{2n}$ such that for any $y\in\mathbb{R}^{n}$, there exists $u_{1},u_{2},...u_{2n}\in\mathbb{R}^{d}$ satisfying
$\sum_{i=1}^{2n}\tilde{D}_{i}Xu_{i}=y.$
However, we can further prove that using randomly sampled hyperplane arrangement patterns $\tilde{D}_{i}$, we can fit arbitrary vector $y$ with probability at least $1-\delta$, provided that we sample sufficiently many hyperplane arrangement patterns.
Proposition 2.
Suppose we sampled $\tilde{P}=2\kappa\log(\frac{n}{\delta})$ hyperplane arrangement patterns, provided that $\mathcal{M}$ is invertible. Then, with probability at least $1-\delta$, for any $y\in\mathbb{R}^{n}$ there exists $u_{1},u_{2},...,u_{\tilde{P}}$ satisfying
$\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}=y.$
Proposition 2 shows that when we sample sufficiently many hyperplane arrangement patterns, we can fit arbitrary vector with high probability, hence $\tilde{p}_{1}^{*}=0$. The two propositions lead to the following approximation result.
Corollary 1.
When $\beta=0$ and $\tilde{P}\geq 2\kappa\log(\frac{n}{\delta}),$ $p_{0}^{*}=p_{1}^{*}=0$ with probability at least $1-\delta$.
One natural question is “how large will $\kappa$ be?” Interestingly, when (A1) is satisfied, we can show that $\kappa=O(n/d)$ holds (see Section 3.5 for details). Hence, for random data, we use approximately $O(\log n)$ neurons to fit to $y\in\mathbb{R}^{n}$ with high probability. Note that we need at least $n/d$ hyperplane arrangement patterns and $n$ parameters to fit to any vector. This means that for random data, Proposition 2 is optimal up to a logarithmic factor.
3.2 Unconstrained Relaxation with $l_{2}$ Regularization
In the case of $l_{2}$ regularization, we can find that when we sample more hyperplane arrangement patterns, the optimal value decreases with the same order as the number of planes with high probability. The proof is rather straightforward: we solve the linear-constrained quadratic problem, and then use matrix chernoff bounds to relate the eigenvalues of $\mathcal{M}$ with the eigenvalue of sample mean of matrices
$\mathcal{M}_{\tilde{P}}=\frac{1}{\tilde{P}}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i}.$
Theorem 4.
Let $p_{2}^{*}$ be the optimal value of the unconstrained relaxed problem with $l_{2}$ regularization, and suppose $\tilde{P}\geq 12\kappa\log(\frac{2n}{\delta})$. With probability at least $1-\delta$ there exists scalars $C_{1},C_{2}>0$ that satisfies
$\frac{C_{1}}{\tilde{P}}\leq p_{2}^{*}\leq\frac{C_{2}}{\tilde{P}}.$
An intuitive explanation of Theorem 4 is that when we sample sufficiently many hyperplane arrangement patterns, we have $\mathcal{M}_{\tilde{P}}\approx\mathcal{M}$ and the optimal value of the $l_{2}$ regularized problem concentrates to $\frac{\beta}{\tilde{P}}y^{T}\mathcal{M}^{-1}y$. This characterization means that for $l_{2}$ regularization, we cannot have relative optimality gap that is independent of $P$, the total number of all possible hyperplane arrangement patterns.
3.3 Unconstrained Relaxation With Group $l_{1}$ Regularization
Different from problem the case with $l_{2}$ regularization, when we have group $l_{1}$ regularization, we can have relative error bounds that are irrelevant with the total number of hyperplane arrangement patterns. We first obtain an upper bound of $\tilde{p}_{1}^{*}$ by using surrogate variables $\kappa_{i}$ and solving the $l_{2}$ regularized problem,
$\min_{\kappa_{i}\in\mathbb{R}}\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}(\kappa_{i}\lVert u_{i}\rVert_{2}^{2}+\frac{1}{\kappa_{i}}).$ (6)
We cannot exactly solve problem (6). However, we can use the ansatz $\kappa_{i}=\frac{\kappa}{\tilde{P}}$ and apply matrix concentration to upper bound the optimal value.
Proposition 3.
Let $\tilde{p}_{1}^{*}$ the optimal value of Approx C-GReLU. Suppose we sampled $\tilde{P}\geq 8\kappa\log(\frac{n}{\delta})$ hyperplane arrangement patterns. With probability at least $1-\delta$,
$\tilde{p}_{1}^{*}\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}},$
provided that $\mathcal{M}$ is invertible.
Then, we lower bound $p_{1}^{*}$ using the dual problem of 5. As the dual problem becomes a maximization of the dual objective with respect to a dual variable $\lambda$, choosing $\lambda$ well can give a meaningful lower bound on $p_{1}^{*}$. In the proof we choose a scalar multiple of $y$ to obtain a bound that has a similar scale with $\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}}$.
Proposition 4.
Let $p_{1}^{*}$ be the optimal value of Full C-GReLU. When we write $\mathcal{M}_{i}=D_{i}XX^{T}D_{i}$,
$G\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}\leq p_{1}^{*},$
where
$G=1-\frac{\beta}{2\max_{i\in[P]}\sqrt{y^{T}\mathcal{M}_{i}y}}.$
A direct corollary simplifies the lower bound. Corollary 2 can be used when $\beta$ is sufficiently smaller than $\lVert y\rVert\sqrt{\lambda_{\min}(\mathcal{M})}$. As $\lVert y\rVert$ is the scale of $O(\sqrt{n})$ and $\lambda_{\min}(\mathcal{M})$ grows with $d$ for random data matrix $X$, it is likely that the value $\lVert y\rVert\sqrt{\lambda_{\min}(\mathcal{M})}$ is much larger than $\beta$ for most cases.
Corollary 2.
Suppose further that $\lVert y\rVert\sqrt{\lambda_{\min}(\mathcal{M})}\geq\beta$. Then,
$\frac{\beta}{2}\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}\leq p_{1}^{*}.$
By using Proposition 3 and Proposition 4, we obtain a relative error bound between $p_{1}^{*}$ and $\tilde{p}_{1}^{*}$.
Theorem 5.
Let $p_{1}^{*}$ and $\tilde{p}_{1}^{*}$ be optimal values of problem (5) with all possible hyperplane arrangements and randomly sampled arrangements, respectively. Suppose we sampled $\tilde{P}\geq 8\kappa\log(\frac{n}{\delta})$ hyperplane arrangement patterns and $\mathcal{M}$ is invertible. We have
$0\leq\tilde{p}_{1}^{*}-p_{1}^{*}\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}},$
and
$p_{1}^{*}\leq\tilde{p}_{1}^{*}\leq\frac{\sqrt{2\kappa}}{G}p_{1}^{*},$
with probability at least $1-\delta$.
Note that different from the $l_{2}$ regularized case, in this case we have a relative error bound that is independant of the total number of hyperplane arrangement patterns.
A better approximation is possible if we take into account two things: one is that we can use $\max_{i}\sqrt{y^{T}\mathcal{M}_{i}y}/\lVert y\rVert_{2}$ instead of $\lambda_{\max}(XX^{T})$ to bound the relative error with a tighter bound, and we can also use $\mathcal{M}_{\tilde{P}}$ directly instead of $\lambda_{\min}(\mathcal{M})$. Hence, the overall bound can be improved to
$p_{0}^{*}\leq p_{1}^{*}\leq\frac{\sqrt{2}}{G}\frac{\max_{i}\sqrt{y^{T}\mathcal{M}_{i}y}\sqrt{y^{T}\mathcal{M_{\tilde{P}}}^{-1}y}}{\lVert y\rVert_{2}^{2}}p_{0}^{*},$ (7)
where we sampled sufficiently many hyperplane arrangement patterns to make $\mathcal{M}_{\tilde{P}}$ invertible.
3.4 Connection to the MAX-CUT Problem
There is an interesting connection between the upper bound of the relative error and the MAX-CUT problem. From bound (7), we can bound
$\max_{i}\sqrt{y^{T}\mathcal{M}_{i}y}\leq\sqrt{\max_{b\in\{0,1\}^{n}}b^{T}diag(y)XX^{T}diag(y)b},$
where solving the right hand side is equivalent to solving the max-cut problem of the graph with adjacency matrix
$\frac{1}{4}\begin{bmatrix}I\\ 1^{T}\end{bmatrix}diag(y)XX^{T}diag(y)\begin{bmatrix}I&1\end{bmatrix}.$
Hence, in special cases where $X$ is orthogonally separable or when $X$ has negative correlation, we may further bound the relative error. The connection with MAX-CUT shows that the relative error, or the possibility of approximation, in intrinsically related to the structure of the dataset and its clusterization properties.
3.5 Scale of $\kappa$ for Random Data
In this subsection we show that under (A1), we can show that $\kappa\leq O(c)$ holds with high probability for sufficiently large $d$. This means that for random data, as long as $d$ grows with $n$, we only need to sample $O(\log n)$ hyperplane arrangement patterns to obtain a constant factor approximation of the unconstrained relaxation. This bound is more practical compared to most bounds that require networks of width at least $n$.
To prove the identity, we first use the non-asymptotic inequality deduced from Gordon’s minimax comparison
$\lambda_{max}(XX^{T})\leq 2(\sqrt{n}+\sqrt{d})^{2},$
which holds with high probability. To upper bound $\kappa$, finding a lower bound of $\lambda_{\min}(\mathcal{M})$ is enough.
In order to lower bound $\lambda_{\min}(\mathcal{M})$, we find a closed form expression of $\mathcal{M}_{ij}$ and use Lemma 6 to approximate $\mathcal{M}$ via linearization. We need $d$ to be sufficiently large in order to apply concentration inequalities on $\lVert x_{i}\rVert_{2}$, the norm of rows of $X$.
Theorem 6.
Assume (A1) and suppose $c\geq 1,\delta>0$. There exists $d_{1}$ such that if $d\geq d_{1}$, with probability at least $1-\delta-\frac{1}{(2n)^{8}}$, we have
$\lambda_{min}(\mathcal{M})\geq\frac{d}{10}.$
Hence, $\kappa$ is indeed upper bounded with a constant factor when $d$ is sufficiently large. Using the fact directly leads to the following corollary.
Corollary 3.
When the conditions of Theorem 6 holds, with probability at least $1-\delta-\frac{1}{(2n)^{8}}$, we have
$\kappa\leq 10\sqrt{2}(\sqrt{c}+1)^{2}.$
Moreover, when we sample $\tilde{P}\geq 160(\sqrt{c}+1)^{2}\log(\frac{n}{\delta^{\prime}})$ hyperplane arrangement patterns, we have
$p_{1}^{*}\leq\tilde{p}_{1}^{*}\leq\frac{2\sqrt{10}}{G}(\sqrt{c}+1)p_{1}^{*}$
with probability at least $1-\delta-\delta^{\prime}-\frac{1}{(2n)^{8}}.$
4 Extension to the Constrained Problem
In this section, we move a step further by discussing the relative error between the optimal value of the convex reformulation and its random relaxation. Let ${w_{i}^{*}}_{i=1}^{m}$ be the solution of the unconstrained relaxation of the original problem and let $u_{i}^{*},v_{i}^{*}\in\mathcal{K}_{D_{i}}$ satisfying $w_{i}^{*}=u_{i}^{*}-v_{i}^{*}$ with minimal $\lVert u_{i}^{*}\rVert_{2}+\lVert v_{i}^{*}\rVert_{2}$.
Although there is no universal constant $C$ that satisfies
$\lVert u_{i}^{*}\rVert_{2}+\lVert v_{i}^{*}\rVert_{2}\leq C\lVert w_{i}^{*}\rVert_{2},$
for any cone $\mathcal{K}$ [34], we can show that for random data $X$, it is likely that there exists reasonably large $\mathcal{C}$ that satisfies
$\lVert u_{i}^{*}\rVert_{2}+\lVert v_{i}^{*}\rVert_{2}\leq\mathcal{C}\lVert w_{i}^{*}\rVert_{2},$
for all $\mathcal{K}_{\tilde{D}_{i}}$ and $w_{i}^{*}$ with high probability. In 2 dimensions, $\mathcal{C}$ is directly related with the angle between two rays of the convex cone make. Hence, $\mathcal{C}$ is analogous to the ‘sharpness of each cone $\mathcal{K}_{\tilde{D}_{i}}$’. We defer the proofs to Appendix C.
4.1 Cone sharpness $\mathcal{C}$
We start from defining the cone sharpness constant $\mathcal{C}(\mathcal{K},z)$ for a given convex cone $\mathcal{K}$.
Definition 1.
For a cone $\mathcal{K}$ and unit vector $z$, the sharpness with respect to $z$ is defined as
$C(\mathcal{K},z)=\min_{u,v\in\mathcal{K},u-v=z}\lVert u\rVert_{2}+\lVert v\rVert_{2}.$
We can upper bound the unconstrained relaxation with the constrained relaxation using the cone sharpness.
Proposition 5.
Let $\tilde{p}_{0}^{*}$ and $\tilde{p}_{1}^{*}$ be optimal values of the Gaussian relaxation of the convex reformulation and its unconstrained version, respectively. Suppose the unconstrained problem has solutions $w_{i}^{*}$ for $i\in[\tilde{P}]$, and let
$C=\max_{i\in[\tilde{P}]}\mathcal{C}(\mathcal{K}_{\tilde{D}_{i}},\frac{w_{i}^{*}}{\lVert w_{i}^{*}\rVert_{2}}).$
Then, $\tilde{p}_{0}^{*}\leq C\tilde{p}_{1}^{*}$ holds.
Now, we introduce a strategy to upper bound $\mathcal{C}(\mathcal{K}_{\tilde{D}_{i}},z)$. An interesting fact is that for when (A1) is satisfied, the cone sharpness constant is not large with high probability, and can be bounded by $O(\sqrt{\log n})$ factor. The intuition here is that when $n\asymp d$, we have approximately $O((n/d)^{d})$ many cones - which is only exponential in $d$ when $n/d=c$. Hence, in $d$ dimensions, the number of cone constraints are not that many, and we can have reasonable upper bounds on the sharpness.
We first start with a simple proposition that upper bounds $C(\mathcal{K}_{\tilde{D}_{i}},z)$. The idea is similar to Chebyshev centers of a polyhedron [35].
Proposition 6.
Take any unit vector $z$, and suppose $\mathcal{K}=\{u|(2D-I)Xu\geq 0\}$. If there exists vector $u$ that satisfies
$\lVert u\rVert_{2}\leq 1,\quad(2D-I)Xu\geq\epsilon\cdot|(2D-I)Xz|,$
we know that
$C(\mathcal{K},z)\leq 1+\frac{1}{\epsilon}.$
When (A1) is satisfied, with the rotational invariance of $X$, we can rotate the cone to contain $e_{1}$ and the distribution of other elements on $X$ will not change. Therefore, without loss of generality, we may assume that $\mathcal{K}=\{u\ |\ \tilde{X}u\geq 0\},$ where the first column of $\tilde{X}$ is the absolute value sampled from $\mathcal{N}(0,1)$ and the other columns are sampled from $\mathcal{N}(0,1)$.
What we wish to do is to construct a vector $u$ that satisfies the conditions in Proposition 6. We find such vector using a novel property of $n\times n$ Gaussian random matrices. An application of Gordon’s minimax theorem leads to the following proposition.
Lemma 2.
Suppose $X$ is a $d\times d$ random matrix with entries i.i.d. and sampled from $\mathcal{N}(0,1)$. Also, suppose $b$ is an arbitrary vector that satisfies $\lVert b\rVert_{2}\leq 2\sqrt{n}$ and $n/d=c$. Then, for some $B>0$, with probability at least $1-8\exp(-Bn)$, there exists a vector that satisfies $\lVert\lambda\rVert_{2}\leq 5c$ and $X\lambda\geq b$.
By using the $\lambda$ in Lemma 2, we can construct $u$ for the rotated cone $\mathcal{K}=\{u\ |\ \tilde{X}u\geq 0\}$. Eventually, we can bound $\mathcal{C}(\mathcal{K},z)$ with a logarithmic factor.
Theorem 7.
Let $p_{c}$ to be the probability that the $0.75d-th$ absolute value of $n$ samples sampled uniformly at random from $\mathcal{N}(0,1)$ is less than $d/2n=1/2c$. Also, suppose $d\geq d_{2}$, where $d_{2}$ is the threshold that we can guarantee the existence of such $\lambda\in\mathbb{R}^{d_{2}}$ from Lemma 2 with probability at least $1-\delta^{\prime\prime}$.
Then, with probability at least $1-p_{c}-2/(2n)^{9}-\delta^{\prime\prime}$, we know that
$C(\mathcal{K}_{\tilde{D}_{i}},z)\leq 1+80c^{2}\sqrt{\log(2n)}.$
A direct corollary is that we can approximate the convex reformulation with its unconstrained version, having $O(\sqrt{\log n})$ scale relative bound (Corollary 7).
4.2 Proof of the main results
After the bound on the cone sharpness, the proof of the main theorems follow almost immediately. By considering the convex reformulation of the original problem (1), and first upper bounding it with the unconstrained problem, then upper bounding it again with the unconstrained problem with random relaxation, we can find relative optimality gap between the original problem and its convex relaxation. It is clear that we immediately get a tractable polynomial-time randomized algorithm by solving the randomly relaxed convex problem. At last, by identifying stationary points of the original problem as global minimum of randomly subsampled convex problems, we obtain Theorem 3. See Appendix D for a detailed proof.
5 Conclusion
In this paper, we provided guarantees of approximating the equivalent convex program given in [12] with a much smaller random subprogram. With assumptions on $X$ and the dimension $d$, we proved that the optimal value of the subsampled convex program approximates the full convex program up to a logarithmic factor. Using the approximation results we discuss novel insights on training two-layer neural networks, by showing that under mild assumptions local gradient methods converge to stationary points with optimality guarantees, and propose a practical algorithm to train neural networks with guarantees.
We hope to improve the work in two ways: First, removing the logarithmic factor of the approximation would be an important problem to tackle. Also, extending the theorems to different architectures, i.e. CNNs, transformers, multi-layer networks would be meaningful.
6 Acknowledgements
This work was supported in part by the National Science Foundation (NSF) under Grant DMS-2134248; in part by the NSF CAREER Award under Grant CCF-2236829; in part by the U.S. Army Research Office Early Career Award under Grant W911NF-21-1-0242; in part by the Precourt Institute for Energy and the SystemX Alliance at Stanford University.
Appendix A Useful Lemmas
Lemma 3.
(Matrix Chernoff [36], [37] ) Let $X_{1},X_{2},...,X_{k}$ be independent, random, symmetric real $n\times n$ matrix with $0\preccurlyeq X_{i}\preccurlyeq RI$. Let $\mu_{min}I\preccurlyeq\sum_{i=1}^{k}\mathbb{E}[X_{i}]\preccurlyeq\mu_{max}I$. Then, for all $\delta\in[0,1]$,
$\mathbb{P}(\lambda_{max}(\sum_{i=1}^{k}X_{i})\geq(1+\delta)\mu_{max})\leq ne^{-\delta^{2}\mu_{max}/3R}$
and
$\mathbb{P}(\lambda_{min}(\sum_{i=1}^{k}X_{i})\leq(1-\delta)\mu_{min})\leq ne^{-\delta^{2}\mu_{min}/2R}$
holds.
Lemma 4.
(Gordon’s minimax theorem [38]) Let $\Phi(X)$ be
$\Phi(X)=\min_{x\in S_{x}}\max_{y\in S_{y}}y^{T}Xx+\psi(x,y),$
where $X$ is a random matrix with i.i.d. Gaussian entries, and $\phi(g,h)$ be
$\phi(g,h)=\min_{x\in S_{x}}\max_{y\in S_{y}}\lVert x\rVert_{2}g^{T}y+\lVert y\rVert_{2}h^{T}x+\psi(x,y),$
where entries of $g,h$ are i.i.d. and sampled from $\mathcal{N}(0,1)$. Then, we have
$\mathbb{P}(\Phi(X)\leq M)\leq 2\mathbb{P}(\phi(g,h)\leq M).$
Lemma 5.
(Weyl’s theorem on eigenvalues) Suppose $A,B\in\mathbb{R}^{n\times n}$ are symmetric. Then, the below estimates on the minimum eigenvalue of $A,B$ hold.
$\lambda_{min}(A)\geq\lambda_{min}(B)-\lVert A-B\rVert_{F},$ (8)
$\lambda_{min}(A+B)\geq\lambda_{min}(A)+\lambda_{min}(B).$ (9)
Here, $\rho(A)$ denotes the spectral radius of $A$, and $\lVert A\rVert_{F}$ is the Frobenius norm of $A$.
Lemma 6.
(Spectrum of kernel random matrices [39]) Suppose we obtain $n$ i.i.d. vectors $x_{i}$ from $\mathcal{N}(0,I_{d})$. Let’s consider the kernel matrix
$K_{i,j}=f(\frac{x_{i}\cdot x_{j}}{d}).$
We assume that:
(a) $n/d,d/n$ are bounded as $d\rightarrow\infty$.
(b) $f$ is a $C^{1}$ function in a neighborhood of 1, and a $C^{3}$ function in a neighborhood of 0.
Under the assumptions, the kernel matrix $M$ can (in probability) be approximated consistently in operator norm, when $d,n\rightarrow\infty$, by the matrix $K^{\prime}$, where
$K^{\prime}=(f(0)+\frac{f^{\prime\prime}(0)}{2d})11^{T}+\frac{f^{\prime}(0)}{d}XX^{T}+(f(1)-f(0)-f^{\prime}(0))I_{n}.$
Appendix B Proofs in Section 3.
Proposition 7.
(Proposition 3.1. of the paper) Suppose no two rows of $X$ are parallel. There exists hyperplane patterns $\tilde{D}_{1}$, $\tilde{D}_{2}$, …, $\tilde{D}_{2n}$ such that for all $y\in\mathbb{R}^{n}$, there exists $u_{1},u_{2},...u_{2n}\in\mathbb{R}^{d}$ satisfying
$\sum_{i=1}^{2n}\tilde{D}_{i}Xu_{i}=y.$
Proof.
Let’s begin with a simple observation: for any $i\in[n]$, there exists two different hyperplane pattern $D_{1,i}$ and $D_{2,i}$ satisfying
$|D_{1,i}-D_{2,i}|=diag(e_{i}).$
The construction is intuitive when thought geometrically, passing through only the $i$th plane. A rigorous construction is as below: First, choose a point $c\in\mathbb{R}^{d}$ satisfying
$c\cdot X_{i}=0,\quad c\cdot X_{j}\neq 0\quad for\ all\quad j\neq i,$
where $X_{1},X_{2},...,X_{n}$ are rows of $X$. We can choose such point because the set $\{u|u\cdot X_{i}=0\}\cap\{u|u\cdot X_{j}=0\}$ has measure 0 in $\{u|u\cdot X_{i}=0\}$, provided no two rows of $X$ are parellel. Now, we choose $\epsilon$ to be small such that
$\epsilon=\min_{j\neq i,X_{j}\cdot X_{i}\neq 0}\frac{|c\cdot X_{j}|}{2|X_{i}\cdot X_{j}|}\ >0.$
and take two points $c+\epsilon X_{i}$, $c-\epsilon X_{i}$. Now, for the $i$th hyperplane,
$(c+\epsilon X_{i})\cdot X_{i}>0,\quad(c-\epsilon X_{i})\cdot X_{i}<0,$
and for all other hyperplanes $j\neq i$,
$\epsilon|X_{i}\cdot X_{j}|\leq\frac{1}{2}|c\cdot X_{j}|,$
thus the sign of $(c+\epsilon X_{i})\cdot X_{j},c\cdot X_{j},(c-\epsilon X_{i})\cdot X_{j}$ are identical. This means the hyperplane patterns of $c+\epsilon X_{i}$ and $c-\epsilon X_{i}$ differ only for the $i$th plane, meaning we have found two planes $D_{1,i},D_{2,i}$ such that $|D_{1,i}-D_{2,i}|=diag(e_{i}).$ Now, define $\tilde{D}_{2i-1}=D_{1,i},\tilde{D}_{2i}=D_{2,i}$ for $i\in[n]$, and choose $u_{i}$ vectors such that
$(Xu_{i})[i,1]=y[i,1]$
where the notation $A[i,j]$ denotes the $i$th row, $j$th column of $A$. We can always choose such a vector, as we can choose arbitrary vector $v$ and scale $c_{i}$ until it matches $c_{i}(Xv)[i,1]=y[i,1]$. Hence, we have found such $\tilde{D}$s that we proposed earlier. ∎
Proposition 8.
(Proposition 3.2. of the paper) Suppose we sampled
$\tilde{P}=2\kappa\log(\frac{n}{\delta})$
hyperplane arrangement patterns, provided that $\mathcal{M}$ is invertible. Then, with probability at least $1-\delta$, there exists $u_{1},u_{2},...,u_{\tilde{P}}$ satisfying
$\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}=y.$
Proof.
Think of $X_{i}=\tilde{D}_{i}XX^{T}\tilde{D}_{i}$ as a random independent series of symmetric real $n\times n$ matrices. Also, we know that for any vector $u\in\mathbb{R}^{n}$,
$\sqrt{u^{T}X_{i}u}\leq\lambda_{max}(XX^{T})\lVert\tilde{D}_{i}u\rVert_{2}\leq\lambda_{max}(XX^{T})\lVert u\rVert_{2},$
hence we know that for all $X_{i}$s, $X_{i}\preccurlyeq\lambda_{max}(XX^{T})I$. Now, from Lemma 3, take $\delta=1$. Then, we obtain
$\mathbb{P}(\lambda_{min}(\sum_{i=1}^{\tilde{P}}X_{i})\leq 0)\leq ne^{-\mu_{min}/2\lambda_{max}(XX^{T})}$
where $\mu_{min}=\tilde{P}\lambda_{min}(\mathcal{M})$. Hence, when we plug in
$\tilde{P}=2\kappa\log(\frac{n}{\delta}),$
we can see that $\mathbb{P}(\lambda_{min}(\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})\leq 0)\leq\delta$, and $\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i}$ is invertible with probability $1-\delta$.
At last, consider the augmented matrix $\mathcal{X}=[\tilde{D}_{1}X|\tilde{D}_{2}X|...|\tilde{D}_{\tilde{P}}X]$, which is a $n\times\tilde{P}d$ matrix. As $\mathcal{X}\mathcal{X}^{T}=\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i}$ is invertible, we can see that $\mathcal{X}$ has $n$ nonzero singular values, hence has rank $n$. This means that the column space of $[\tilde{D}_{1}X|\tilde{D}_{2}X|...|\tilde{D}_{\tilde{P}}X]$ also has rank $n$, and we can find vectors $u_{1},u_{2},...,u_{\tilde{P}}$ that satisfies
$\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}=y$
for any given $y\in\mathbb{R}^{n}$. ∎
The two propositions directly lead to the following corollary.
Corollary 4.
(Corollary 3.3. of the paper) Suppose $p_{0}^{*}$ and $p_{1}^{*}$ are solutions to the unconstrained problem that uses all possible hyperplane arrangement patterns and its randomized relaxation, respectively, and the regularization $\beta=0$. Also, suppose $\tilde{P}\geq 2\kappa\log(\frac{n}{\delta})$. Then, $p_{0}^{*}=p_{1}^{*}=0$ with probability at least $1-\delta$.
Proof.
From Proposition 1, we can find $v_{1},v_{2},...,v_{2n}$ that satisfies
$\sum_{i=1}^{2n}D^{\prime}_{i}Xv_{i}=y$
for specific $D^{\prime}_{i}$s. Hence, for the unconstrained problem with all possible hyperplane arrangements, we can choose $u_{i}$s to be $u_{i}=v_{i}$ if $D_{i}=D^{\prime}_{i}$, $u_{i}=0$ otherwise to perfectly fit a given vector $y\in\mathbb{R}^{n}$. Also, from Proposition 2, we can fit a given vector $y$ with probabillity at least $1-\delta$ when we sample hyperplane arrangement patterns more than $2\kappa\log(\frac{n}{\delta})$ times. Hence, with probability at least $1-\delta$, we know that $p_{1}^{*}=0$. ∎
Theorem 8.
(Theorem 3.4. of the paper) Suppose
$p^{*}=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}^{2}.$ (10)
Then, with probability at least $1-\delta$ there exists scalars $C_{1},C_{2}>0$ that satisfies
$\frac{C_{1}}{\tilde{P}}\leq p^{*}\leq\frac{C_{2}}{\tilde{P}},$
provided that $\tilde{P}\geq 12\kappa\log(\frac{2n}{\delta})$ and $\mathcal{M}$ is invertible.
Proof.
Solving the $L_{2}$ regularized problem (10) is equivalent to solving
$\min_{u_{i}\in\mathbb{R}^{d},w}\frac{1}{2}\lVert w-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}^{2}.$ (11)
subject to
$w=\sum_{i=1}^{\tilde{P}}D_{i}Xu_{i}.$
The lagrangian of problem (11) becomes
$L(u,w,\lambda)=\frac{1}{2}\lVert w-y\rVert_{2}^{2}+\lambda^{T}(w-\sum_{i=1}^{\tilde{P}}D_{i}Xu_{i})+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}^{2}.$
As the constraint is linear equality and the objective is convex with respect to the arguments, strong duality holds and
$p^{*}=\min_{u_{i}\in\mathbb{R}^{d},w\in\mathbb{R}^{n}}\max_{\lambda\in\mathbb{R}^{n}}L(u_{i},w,\lambda)=\max_{\lambda\in\mathbb{R}^{n}}\min_{u_{i}\in\mathbb{R}^{d},w\in\mathbb{R}^{n}}L(u_{i},w,\lambda).$
For a given $\lambda$, we can optimize for $u_{i}$ and $w$ to minimize $L(u_{i},w,\lambda)$, where we get $w=y-\lambda$ and $u_{i}=\frac{1}{\beta}X^{T}\tilde{D}_{i}\lambda$. Substituting leads to
$p^{*}=\max_{\lambda\in\mathbb{R}^{n}}-\frac{1}{2}\lVert\lambda\rVert_{2}^{2}+\lambda^{T}y-\frac{1}{2\beta}\lambda^{T}\big{(}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i}\big{)}\lambda.$
Now, let’s write $\frac{1}{\tilde{P}}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i}=\mathcal{M}_{\tilde{P}}$. Then, the dual problem can be simplified as
$p^{*}=\max_{\lambda\in\mathbb{R}^{n}}-\frac{1}{2\beta}\lambda^{T}(\beta I+\tilde{P}\mathcal{M}_{\tilde{P}})\lambda+\lambda^{T}y$
and the optimum $p^{*}$ is given as
$p^{*}=\frac{\beta}{2}y^{T}(\beta I+\tilde{P}\mathcal{M}_{\tilde{P}})^{-1}y.$
When $\tilde{P}\geq 12\kappa\log(\frac{2n}{\delta})$, by Lemma 3 we can see that
$\mathbb{P}\bigg{(}\lambda_{min}(\mathcal{M}_{\tilde{P}})\geq\frac{\lambda_{min}(\mathcal{M})}{2}\bigg{)}\geq 1-\frac{\delta}{2}$
and
$\mathbb{P}\bigg{(}\lambda_{max}(\mathcal{M}_{\tilde{P}})\leq\frac{3\lambda_{max}(\mathcal{M})}{2}\bigg{)}\geq 1-\frac{\delta}{2}.$
Hence, with probability at least $1-\delta$, we can see that
$(\beta+\frac{\tilde{P}}{2}\lambda_{\min}(\mathcal{M}))I\preccurlyeq\tilde{P}\mathcal{M}_{\tilde{P}}+\beta I\preccurlyeq(\beta+\frac{3\tilde{P}}{2}\lambda_{\max}(\mathcal{M}))I$
and we can find $\mathcal{R}_{1},\mathcal{R}_{2}>0$ that satisfies
$\mathcal{R}_{1}\tilde{P}I\preccurlyeq\tilde{P}\mathcal{M}_{\tilde{P}}+\beta I\preccurlyeq\mathcal{R}_{2}\tilde{P}I.$
An example is $\mathcal{R}_{1}=\frac{\lambda_{\min}(\mathcal{M})}{2}$, $\mathcal{R}_{2}=\frac{3\lambda_{\min}(\mathcal{M})}{2}+1$, provided that $\tilde{P}$ is larger than $\beta$. This means that
$\frac{\beta\lVert y\rVert_{2}^{2}}{2\mathcal{R}_{2}\tilde{P}}\leq p^{*}\leq\frac{\beta\lVert y\rVert_{2}^{2}}{2\mathcal{R}_{1}\tilde{P}},$
and take $C_{1}=\frac{\beta\lVert y\rVert_{2}^{2}}{2\mathcal{R}_{2}},C_{2}=\frac{\beta\lVert y\rVert_{2}^{2}}{2\mathcal{R}_{1}}$ to finish the proof. ∎
Proposition 9.
(Proposition 3.5. of the paper) Suppose
$p_{1}^{*}=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2},$
and $\tilde{P}\geq 8\kappa\log(\frac{n}{\delta})$. Then with probability at least $1-\delta$,
$p_{1}^{*}\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}},$
provided that $\mathcal{M}$ is invertible.
Proof.
We know that
$p_{1}^{*}$ $=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}$
$=\min_{\kappa_{i}\in\mathbb{R},u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}(\kappa_{i}\lVert u_{i}\rVert_{2}^{2}+\frac{1}{\kappa_{i}})$
$=\min_{\kappa_{i}\in\mathbb{R}}\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}(\kappa_{i}\lVert u_{i}\rVert_{2}^{2}+\frac{1}{\kappa_{i}}).$
With the same idea from the proof of Equation 10, we can see that when $\kappa_{i}$s are given, the dual problem becomes
$\max_{\lambda\in\mathbb{R}^{n}}-\frac{1}{2\beta}\lambda^{T}(\beta I+\sum_{i=1}^{\tilde{P}}\frac{1}{\kappa_{i}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})\lambda+\lambda y,$
and the inner minimization problem has the minimum
$\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}(\kappa_{i}\lVert u_{i}\rVert_{2}^{2}+\frac{1}{\kappa_{i}})=\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}\frac{1}{\kappa_{i}}+\frac{\beta}{2}y^{T}(\beta I+\sum_{i=1}^{\tilde{P}}\frac{1}{\kappa_{i}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})^{-1}y.$
Hence, we can see that
$p_{1}^{*}$ $=\min_{\kappa_{i}\in\mathbb{R}}\frac{\beta}{2}\sum_{i=1}^{\tilde{P}}\frac{1}{\kappa_{i}}+\frac{\beta}{2}y^{T}(\beta I+\sum_{i=1}^{\tilde{P}}\frac{1}{\kappa_{i}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})^{-1}y$
$\leq\frac{\beta}{2}\min_{\kappa\in\mathbb{R}}y^{T}(\beta I+\frac{\kappa}{\tilde{P}}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})^{-1}y+\kappa$
$\leq\frac{\beta}{2}\min_{\kappa\in\mathbb{R}}\frac{y^{T}(\frac{1}{\tilde{P}}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})^{-1}y}{\kappa}+\kappa$
$\leq\beta\sqrt{y^{T}(\frac{1}{\tilde{P}}\sum_{i=1}^{\tilde{P}}D_{i}XX^{T}D_{i})^{-1}y}$
$\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\lambda_{\min}(\mathcal{M})}$
Where the last inequality follows from matrix Chernoff that with probability at least $1-\delta$, the minimum eigenvalue
$\lambda_{\min}(\frac{1}{\tilde{P}}\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}XX^{T}\tilde{D}_{i})\geq\frac{\lambda_{\min}(\mathcal{M})}{2}.$
Proposition 10.
(Proposition 3.6. of the paper) Suppose
$p_{0}^{*}=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{P}D_{i}Xu_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{P}\lVert u_{i}\rVert_{2}.$
Here $D_{i}$s are all possible hyperplane arrangement patterns. Also, write $\mathcal{M}_{i}=D_{i}XX^{T}D_{i}$. Then,
$G\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}\leq p_{0}^{*},$
where
$G=1-\frac{\beta}{2\max_{i\in[P]}\sqrt{y^{T}\mathcal{M}_{i}y}}$
Proof.
Let’s think of the dual problem of the original optimization problem. The lagrangian $L(u_{i},w,\lambda)$ is given as
$L(u_{i},w,\lambda)=\frac{1}{2}\lVert w-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}+\lambda^{T}(w-\sum_{i=1}^{\tilde{P}}D_{i}Xu_{i}),$
and the dual problem is $\max_{\lambda}\min_{u_{i},w}L(u_{i},w,\lambda)$. As there is only a linear equality constraint and strong duality holds, we can see that
$p_{0}^{*}=\max_{\lambda}\min_{u_{i},w}\frac{1}{2}\lVert w-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert u_{i}\rVert_{2}+\lambda^{T}(w-\sum_{i=1}^{\tilde{P}}D_{i}Xu_{i}).$
When $\lambda$ is fixed, the inner minimization problem can be solved as:
i) If $\lVert X^{T}D_{i}\lambda\rVert_{2}>\beta$ for some $i$: Take $u_{i}=CX^{T}D_{i}\lambda$ and send $C$ to infinity to obtain $\infty$- ∞ as the solution to the inner minimization problem.
ii) If $\lVert X^{T}D_{i}\lambda\rVert_{2}\leq\beta$ for all $i$: Each $u_{i}=0$ when the inner minimization problem is solved, due to Cauchy-Schwartz and
$\beta\lVert u_{i}\rVert_{2}\geq\lVert X^{T}D_{i}\lambda\rVert_{2}\lVert u_{i}\rVert_{2}\geq\lambda^{T}D_{i}Xu_{i}$
holds for all $u_{i}\in\mathbb{R}^{d}$. Also, $w=y-\lambda$ should hold, and the objective becomes maximizing $-\frac{1}{2}\lVert\lambda\rVert_{2}^{2}+\lambda^{T}y$.
From i), ii), we can see that the dual problem becomes
$\max_{\lambda}-\frac{1}{2}\lVert\lambda\rVert_{2}^{2}+\lambda^{T}y$
subject to
$\lVert X^{T}D_{i}\lambda\rVert_{2}\leq\beta\quad\forall i\in[P].$
Now, let’s find the maximal scaling coefficient $k$ and the corresponding $\lambda_{k}=ky$ that meets all the constraints. For that $\lambda$, it is clear that $p_{0}^{*}\geq-\frac{1}{2}\lVert\lambda_{k}\rVert_{2}^{2}+\lambda_{k}^{T}y$. When we substitute $ky$ to $\lambda$, we get the constraint of $k$ for each $i$:
$k\lVert X^{T}D_{i}y\rVert_{2}\leq\beta,\quad k\leq\frac{\beta}{\sqrt{y^{T}D_{i}XX^{T}D_{i}y}}$
must hold. When we write $D_{i}XX^{T}D_{i}=\mathcal{M}_{i}$, $k$ should satisfy
$k\leq\frac{\beta}{\sqrt{y^{T}\mathcal{M}_{i}y}}$
for all $i\in[\tilde{P}]$. Choose $k=\frac{\beta}{\max_{i}\sqrt{y^{T}\mathcal{M}_{i}y}}$. Substituting $ky$ to the dual problem leads
$p_{0}^{*}\geq-\frac{\beta^{2}\lVert y\rVert_{2}^{2}}{2\max_{i}\{y^{T}\mathcal{M}_{i}y\}}+\frac{\beta\lVert y\rVert_{2}^{2}}{\max_{i}\{\sqrt{y^{T}\mathcal{M}_{i}y}\}}=G\frac{\beta\lVert y\rVert_{2}^{2}}{\max_{i}\{\sqrt{y^{T}\mathcal{M}_{i}y}\}}.$
At last, we know that $\mathcal{M}_{i}\preccurlyeq\lambda_{\max}(XX^{T})I$ for all $i\in[P]$. Hence, $y^{T}\mathcal{M}_{i}y\leq\lambda_{\max}(XX^{T})\lVert y\rVert_{2}^{2}$ for all $i\in[P]$ and
$p_{0}^{*}\geq G\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}.$
Corollary 5.
(Corollary 3.7. of the paper) Suppose further that $\lVert y\rVert_{2}\sqrt{\lambda_{\min}(\mathcal{M})}\geq\beta$. Then,
$\frac{\beta}{2}\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}\leq p_{1}^{*}.$
Proof.
We know that $G=1-\frac{\beta}{2\max_{i\in[P]}\sqrt{y^{T}\mathcal{M}_{i}y}}\geq 1-\frac{\beta}{2\lVert y\rVert_{2}\sqrt{\lambda_{\min}(\mathcal{M})}}\geq\frac{1}{2}.$
Theorem 9.
(Theorem 3.8. of the paper) Suppose $p_{0}^{*}$ and $p_{1}^{*}$ are solutions to the unconstrained problem that uses all possible hyperplane arrangement patterns and its randomized relaxation, respectively. Suppose further that $\tilde{P}\geq 8\kappa\log(\frac{n}{\delta})$ and $\mathcal{M}$ is invertible. We have
$0\leq p_{1}^{*}-p_{0}^{*}\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}}$
and
$p_{0}^{*}\leq p_{1}^{*}\leq\frac{\sqrt{2\kappa}}{G}p_{0}^{*}$
with probability at least $1-\delta$.
Proof.
The two inequalities directly follow from Proposition 3 and Proposition 4
Proposition 11.
$\mathcal{M}_{ij}=\{\frac{1}{2}-\frac{1}{2\pi}\arccos(\frac{x_{i}\cdot x_{j}}{\lVert x_{i}\rVert_{2}\lVert x_{j}\rVert_{2}})\}(x_{i}\cdot x_{j})$ for all $i,j\in[n]$. Here, $x\cdot y$ denotes the innter product between $x$ and $y$.
Proof.
First, recall that $\mathcal{M}=\mathbb{E}_{g\sim N(0,I_{d})}[diag[\mathbbm{1}(Xg\geq 0)]XX^{T}diag[\mathbbm{1}(Xg\geq 0)]]$. Hence, $\mathcal{M}_{ij}$ has the expression
$\mathcal{M}_{ij}=\mathbb{P}(x_{i}\cdot g\geq 0)\mathbb{P}(x_{j}\cdot g\geq 0)(x_{i}\cdot x_{j}).$
We know that for fixed $x_{i},x_{j}$, the orthant probability is given as
$\mathbb{P}(x_{i}\cdot g\geq 0)\mathbb{P}(x_{j}\cdot g\geq 0)=\frac{1}{2}-\frac{1}{2\pi}\arccos(\frac{x_{i}\cdot x_{j}}{\lVert x_{i}\rVert_{2}\lVert x_{j}\rVert_{2}}),$
where $g\sim\mathcal{N}(0,I_{d})$. This directly implies the claim. ∎
Theorem 10.
(Theorem 3.9. of the paper) Let each row $x_{i}$ of $X$ is sampled i.i.d. from $N(0,I_{d})$. Furthermore, suppose $c,\delta>0$ are given and $n/d=c$ is fixed. There exists $d_{1}$ such that if $d\geq d_{1}$, with probability at least $1-\delta-\frac{1}{(2n)^{8}}$, we have
$\lambda_{min}(\mathcal{M})\geq\frac{d}{10}.$
Proof.
Note that $f(t)=(\frac{1}{2}-\frac{1}{2\pi}\arccos(t))t$ is not differentiable at a neighborhood around $t=1$. Hence, we need $\tilde{f}_{\epsilon}$, a $C^{1}$ approximator of $f$, defined as
$\tilde{f}_{\epsilon}(x)=\begin{cases}f(x)\quad when\quad x\leq 1-\epsilon\\ f^{\prime}(1-\epsilon)x+f(1-\epsilon)-(1-\epsilon)f^{\prime}(1-\epsilon)\quad when\quad x\geq 1-\epsilon.\end{cases}$
Let $\tilde{M}$ be a matrix that satisfies
$[\tilde{M}]_{ij}=\tilde{f}_{\epsilon_{0}}(\frac{x_{i}\cdot x_{j}}{d}),$
where $\epsilon_{0}=\frac{1}{100c}\leq\frac{1}{100}$. From Lemma 6, there exists $d_{2}$ such that for every $d\geq d_{2}$,
$\lambda_{min}(\tilde{M})\geq\lambda_{min}(\tilde{M}^{\prime})-\epsilon_{0}.$
Now, there exists $d_{3}$ such that when $d\geq d_{3}$, we can decompose $\frac{\mathcal{M}}{d}$ as
$\frac{\mathcal{M}}{d}=\tilde{M}+diag(\frac{\lVert x_{i}\rVert_{2}^{2}}{2d}-\tilde{f}_{\epsilon_{0}}(\frac{\lVert x_{i}\rVert_{2}^{2}}{d}))+\mathcal{M}_{e},$
where
$[\mathcal{M}_{e}]_{ij}=\begin{cases}0\quad if\quad i=j\\ \frac{1}{2\pi}(-\arccos({\frac{x_{i}\cdot x_{j}}{\lVert x_{i}\rVert_{2}\lVert x_{j}\rVert_{2}}})+\arccos(\frac{x_{i}\cdot x_{j}}{d}))\frac{x_{i}\cdot x_{j}}{d}\quad if\quad i\neq j.\end{cases}$
with probability at least $21-\delta/$ 2. Such $d_{3}$ exists as for each $i\neq j$, $\frac{x_{i}\cdot x_{j}}{d}\sim\mathcal{N}(0,\frac{1}{d})$ and
$\mathbb{P}(\frac{|x_{i}\cdot x_{j}|}{d}\geq\frac{1}{2})\leq 2\exp(-\frac{d}{8}),$
meaning that for all $i\neq j$,
$\mathbb{P}(\max_{i\neq j}{\frac{|x_{i}\cdot x_{j}|}{d}}\leq\frac{1}{2})\geq 1-2n^{2}\exp(-\frac{d}{8})=1-2c^{2}d^{2}\exp(-\frac{d}{8}),$
and choose $d_{3}$ large enough so that $2c^{2}d^{2}\exp(-\frac{d}{8})\leq\frac{\delta}{2}$. This means with probability at least $21-\delta/$ 2, $f(\frac{x_{i}\cdot x_{j}}{d})=\tilde{f}_{\epsilon_{0}}(\frac{x_{i}\cdot x_{j}}{d})$ for all $i\neq j$, meaning that the off-diagonal entries of $\frac{\mathcal{M}}{d}$ can be decomposed to $\tilde{M}$ and $\mathcal{M}_{e}$. Now, Weyl’s inequality leads to
$\frac{\lambda_{\min}(\mathcal{M})}{d}\geq\lambda_{\min}(\tilde{\mathcal{M}})+\min(\frac{\lVert x_{i}\rVert_{2}^{2}}{2d})-\max(\tilde{f}_{\epsilon_{0}}(\frac{\lVert x_{i}\rVert_{2}^{2}}{d}))-\lVert\mathcal{M}_{e}\rVert_{F}.$ (12)
Then, from the concentration inequality of Chi-square random variables, we know that there exists $d_{4}$ such that for all $d\geq d_{4}$,
$(1-\epsilon_{0})d\leq(1-\frac{\epsilon_{0}}{\log 2n})d\leq\lVert x_{i}\rVert_{2}^{2}\leq(1+\frac{\epsilon_{0}}{\log 2n})d\leq(1+\epsilon_{0})d,\quad\forall\quad i\in[n],$ (13)
with probability at least $21-\delta/$ 2. Hence, when $d\geq\max\{d_{2},d_{3},d_{4}\},$ both inequalities (12) and (13) satisfy with probability at least $1-\delta$. Use (12) and (13) to get
$\frac{\lambda_{\min}(\mathcal{M})}{d}$ $\geq\lambda_{\min}(\tilde{\mathcal{M}})+\min(\frac{\lVert x_{i}\rVert_{2}^{2}}{2d})-\max(\tilde{f}_{\epsilon_{0}}(\frac{\lVert x_{i}\rVert_{2}^{2}}{d}))-\lVert\mathcal{M}_{e}\rVert_{F}$
$\geq\lambda_{\min}(\tilde{\mathcal{M}})+\frac{1}{2}(1-\epsilon_{0})-\tilde{f}_{\epsilon_{0}}(1+\epsilon_{0})-\lVert\mathcal{M}_{e}\rVert_{F}$
$\geq\lambda_{\min}(\tilde{\mathcal{M}^{\prime}})-\epsilon_{0}+\frac{1}{2}(1-\epsilon_{0})-\tilde{f}_{\epsilon_{0}}(1+\epsilon_{0})-\lVert\mathcal{M}_{e}\rVert_{F}.$
The last inequality follows from the fact that $d\geq d_{2}$. We also know that
$\lambda_{\min}(\tilde{M}^{\prime})\geq(\tilde{f}_{\epsilon_{0}}(1)-\tilde{f}_{\epsilon_{0}}(0)-\tilde{f}_{\epsilon_{0}}^{\prime}(0)),$
as both $11^{T},XX^{T}$ are not invertible. Moreover, $\tilde{f}_{\epsilon_{0}}(0)=f(0)=0$, $\tilde{f}_{\epsilon_{0}}^{\prime}(0)=f^{\prime}(0)=1/4$. Substitute to get
$\frac{\lambda_{\min}(\mathcal{M})}{d}$ $\geq\tilde{f}_{\epsilon_{0}}(1)-\frac{1}{4}-\epsilon_{0}+\frac{1}{2}(1-\epsilon_{0})-\tilde{f}_{\epsilon_{0}}(1+\epsilon_{0})-\lVert\mathcal{M}_{e}\rVert_{F}.$
$\geq\frac{1}{4}-\frac{3}{2}\epsilon_{0}-\epsilon_{0}f^{\prime}(1-\epsilon_{0})-\lVert\mathcal{M}_{e}\rVert_{F}$
$\geq 0.209-\lVert\mathcal{M}_{e}\rVert_{F}.$
At last, upper bounding $\lVert\mathcal{M}_{e}\rVert_{F}$ would be enough to prove the claim. First, we know that when $|u|,|v|\leq\frac{2}{3}$, we know that
$|\arccos(u)-\arccos(v)|\leq|\arccos^{\prime}(2/3)||u-v|\leq 2|u-v|.$
This leads to
$\lVert\mathcal{M}_{e}\rVert_{F}$ $\leq\frac{1}{2\pi}\sqrt{\sum_{i=1}^{n}\sum_{j\neq i}|-\arccos({\frac{x_{i}\cdot x_{j}}{\lVert x_{i}\rVert_{2}\lVert x_{j}\rVert_{2}}})+\arccos(\frac{x_{i}\cdot x_{j}}{d})|^{2}|\frac{x_{i}\cdot x_{j}}{d}|^{2}}$
$\leq\frac{2}{2\pi}\sqrt{\sum_{i=1}^{n}\sum_{j\neq i}|-{\frac{x_{i}\cdot x_{j}}{\lVert x_{i}\rVert_{2}\lVert x_{j}\rVert_{2}}}+\frac{x_{i}\cdot x_{j}}{d}|^{2}|\frac{x_{i}\cdot x_{j}}{d}|^{2}}$
$\leq\frac{\epsilon_{0}}{\pi(1-\epsilon_{0})\log(2n)}\sqrt{\sum_{i=1}^{n}\sum_{j\neq i}|\frac{x_{i}\cdot x_{j}}{d}|^{4}}$
with probability at least $1-\delta$. We know from the concentration inequality of maximum of absolute values that
$\mathbb{P}(\max_{i\leq j}|x_{i}\cdot x_{j}|\leq 4\sqrt{2\log(2n)}\lVert x_{i}\rVert_{2})\geq 1-\frac{2}{(2n)^{9}},\quad\forall\quad i\in[n].$ (14)
Hence, with probability at least $1-\delta-\frac{1}{(2n)^{8}},$ we know that both (13) and
$\max_{i\leq j}|x_{i}\cdot x_{j}|\leq 4\sqrt{2\log(2n)\cdot(1+\epsilon_{0})d}$
holds. Hence,
$\lVert\mathcal{M}_{e}\rVert_{F}$ $\leq\frac{\epsilon_{0}}{\pi(1-\epsilon_{0})\log(2n)}\sqrt{\sum_{i=1}^{n}\sum_{j\neq i}|\frac{x_{i}\cdot x_{j}}{d}|^{4}}$
$\leq\frac{\epsilon_{0}}{\pi(1-\epsilon_{0})d^{2}\log(2n)}\sqrt{\sum_{i=1}^{n}\sum_{j\neq i}|x_{i}\cdot x_{j}|^{4}}$
$\leq\frac{\epsilon_{0}n}{\pi(1-\epsilon_{0})d^{2}\log(2n)}\cdot 32\log(2n)(1+\epsilon_{0})d$
$\leq\frac{32\epsilon_{0}(1+\epsilon_{0})c}{\pi(1-\epsilon_{0})}$
$\leq\frac{32}{100\pi}\frac{101}{99}\leq 0.104$
Hence, we know that with probability at least $1-\delta-\frac{1}{(2n)^{8}}$,
$\lambda_{\min}(\mathcal{M})\geq\frac{d}{10}$
holds, provided that $d\geq d_{2},d_{3},d_{4}$. Choose $d_{1}=\max\{d_{2},d_{3},d_{4}\}$. ∎
Corollary 6.
(Corollary 3.10. of the paper) When the conditions of Theorem 6 holds, with probability at least $1-\delta-\frac{1}{(2n)^{8}}$, we have
$\kappa\leq 20(\sqrt{c}+1)^{2}.$
Moreover, when we sample $\tilde{P}\geq 160(\sqrt{c}+1)^{2}\log(\frac{n}{\delta^{\prime}})$ hyperplane arrangement patterns, we have
$p_{1}^{*}\leq\tilde{p}_{1}^{*}\leq\frac{2\sqrt{10}}{G}(\sqrt{c}+1)p_{1}^{*}$
with probability at least $1-\delta-\delta^{\prime}-\frac{1}{(2n)^{8}}.$
Proof.
We know that
$\sqrt{\lambda_{max}(XX^{T})}\leq\sqrt{2}(\sqrt{c}+1)\sqrt{d},$
holds with high probability. Also, from Theorem 6, we know that
$\sqrt{\lambda_{min}(\mathcal{M})}\geq\frac{1}{\sqrt{10}}\sqrt{d}.$
with probability at least $1-\delta-\frac{1}{(2n)^{8}}$. Combine the two results to obtain the wanted upper bound on $\kappa$. Also, we know that with probability at least $1-\delta^{\prime}$,
$p_{1}^{*}\leq\tilde{p}_{1}^{*}\leq\frac{\sqrt{2\kappa}}{G}p_{1}^{*},$
from Theorem 5. Directly using the upper bound on $\kappa$ leads to the following result. ∎
Appendix C Proofs in Section 4.
Proposition 12.
(Proposition 4.2. of the paper) For subsampled hyperplane arrangement patterns $\tilde{D}_{1},\tilde{D}_{2},...,\tilde{D}_{\tilde{P}}$, let
$p_{0}^{*}=\min_{u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}),$ (15)
$p_{1}^{*}=\min_{w_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xw_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert w_{i}\rVert_{2},$ (16)
and $\mathcal{K}_{\tilde{D}_{i}}=\{u|(2\tilde{D}_{i}-I)Xu\geq 0\}$. Suppose problem (16) has solutions $w_{i}^{*}$ for $i\in[\tilde{P}]$, and let
$C=\max_{i\in[\tilde{P}]}C(\mathcal{K}_{\tilde{D}_{i}},\frac{w_{i}^{*}}{\lVert w_{i}^{*}\rVert_{2}}).$
Then, $p_{0}^{*}\leq Cp_{1}^{*}$ holds.
Proof.
For the optimum $w_{i}^{*}$ of problem (16), decompose $w_{i}^{*}$ into $u_{i}^{*},v_{i}^{*}$ such that $\lVert u_{i}^{*}\rVert_{2}+\lVert v_{i}^{*}\rVert_{2}$ is minimal, $u_{i}^{*},v_{i}^{*}\in K_{i}$ and $u_{i}^{*}-v_{i}^{*}=w_{i}^{*}$. Then,
$\frac{\lVert u_{i}^{*}\rVert_{2}+\lVert v_{i}^{*}\rVert_{2}}{\lVert w_{i}^{*}\rVert_{2}}\leq C(\mathcal{K}_{\tilde{D}_{i}},\frac{w_{i}^{*}}{\lVert w_{i}^{*}\rVert_{2}})\leq C$
and when we substitute $u_{i}^{*},v_{i}^{*}$ in (2), we can see that the result would not be greater than $Cp_{1}^{*}$, as the regression loss is identical and the regularization loss does not blow up $C$ times. Hence, we may conclude that $p_{0}^{*}\leq Cp_{1}^{*}$. ∎
Proposition 13.
(Proposition 4.3. in the paper) Take any unit vector $z$ and a cone $\mathcal{K}=\{u|(2D-I)Xu\geq 0\}$. If there exists a vector $u$ that satisfies
$\lVert u\rVert_{2}\leq 1,\quad(2D-I)Xu\geq\epsilon\cdot|(2D-I)Xz|,$
we know that
$C(\mathcal{K},z)\leq 1+\frac{1}{\epsilon}.$
Proof.
As $(2D-I)Xu\geq\epsilon\cdot|(2D-I)Xz|\geq(2D-I)X(\pm\epsilon z),$ we know that the vectors $u+\epsilon z,u-\epsilon z\in\mathcal{K}$. Then, the two vectors
$v_{1}=\frac{1}{2}(\frac{u}{\epsilon}+z),\ \ v_{2}=\frac{1}{2}(\frac{u}{\epsilon}-z).$
become two vectors in $\mathcal{K}$ that satisfy $v_{1}-v_{2}=z$. Hence, the cone sharpness
$C(\mathcal{K},z)\leq\lVert v_{1}\rVert_{2}+\lVert v_{2}\rVert_{2}\leq 1+\frac{1}{\epsilon}\lVert u\rVert_{2}\leq 1+\frac{1}{\epsilon}.$
where the first inequality follows from the definition of $C(\mathcal{K},z)$, the second inequality follows from triangular inequality, and the last follows from the fact that $u$ has norm no greater than 1. ∎
Lemma 7.
(Lemma 4.4. in the paper) Suppose $X$ is a $d\times d$ random matrix with entries i.i.d. and sampled from $\mathcal{N}(0,1)$. Also, suppose $b\in\mathbb{R}^{d}$ is an arbitrary vector that satisfies $\lVert b\rVert_{2}\leq 2\sqrt{n},$ and $n/d=c$. Then, there exists a $B>0$ such that with probability at least $1-8\exp(-Bd)$, there exists a vector that satisfies $\lVert\lambda\rVert_{2}\leq 5c$ and $X\lambda\geq b$.
Proof.
Showing that $\min_{\lambda\in\mathbb{R}^{d}}\ \lVert\lambda\rVert_{2}$ subject to $X\lambda\geq b$ is no larger than $5c$ with high probability is enough to show the theorem. Also, suppose $\lVert b\rVert_{2}\geq\sqrt{n}$, as when we prove for $\lVert b\rVert_{2}\geq\sqrt{n}$, the case when $\lVert b\rVert_{2}\leq\sqrt{n}$ is automatically proved. Denote the optimal $\lVert\lambda\rVert_{2}=p_{0}$. Then, we know that
$p_{0}=\min_{\lambda\in\mathbb{R}^{d}}\max_{\nu\in\mathbb{R}_{-}^{d}}\lVert\lambda\rVert_{2}+\nu^{T}(X\lambda-b),$
by simply writing the Lagrangian of the problem. Hence, when we write
$\Phi_{0}(X)$ $=\min_{\lambda\in\mathbb{R}^{d}}\max_{\nu\in\mathbb{R}_{-}^{d}}\lVert\lambda\rVert_{2}+\nu^{T}(X\lambda-b)$
$=\max_{\nu\in\mathbb{R}_{-}^{d}}\min_{\lambda\in\mathbb{R}^{d}}\lVert\lambda\rVert_{2}+\nu^{T}(X\lambda-b)$
$=-\min_{\nu\in\mathbb{R}_{-}^{d}}\max_{\lambda\in\mathbb{R}^{d}}-\lVert\lambda\rVert_{2}-\nu^{T}(X\lambda-b)$
$=-\min_{\nu\in\mathbb{R}_{+}^{d}}\max_{\lambda\in\mathbb{R}^{d}}-\lVert\lambda\rVert_{2}+\nu^{T}(X\lambda-b),$
and write
$\Phi(X)=\min_{\nu\in\mathbb{R}_{+}^{d}}\max_{\lambda\in\mathbb{R}^{d}}-\lVert\lambda\rVert_{2}+\nu^{T}(X\lambda-b),$
we know that when
$\phi(g,h)=\min_{\nu\in\mathbb{R}_{+}^{d}}\max_{\lambda\in\mathbb{R}^{d}}\lVert\nu\rVert_{2}g^{T}\lambda+\lVert\lambda\rVert_{2}h^{T}\nu-\lVert\lambda\rVert_{2}-\nu^{T}b,$
$\mathbb{P}(\Phi_{0}(X)\leq-M)=\mathbb{P}(\Phi(X)\geq M)=1-\mathbb{P}(\Phi(X)\leq M)\geq 1-2\mathbb{P}(\phi(g,h)\leq M)$
from Gordon’s minimax theorem. Here, $g,h\in\mathbb{R}^{d}$ and entries of $g,h\sim\mathcal{N}(0,1)$.
Hence, finding an almost sure lower bound of $\phi(g,h)$ and negating it gives an almost sure upper bound of $\Phi_{0}(X)$.
Now, let’s solve $\phi(g,h)$. We can rewrite the formula of $\phi(g,h)$ as
$\phi(g,h)=\min_{\nu\in\mathbb{R}_{+}^{d}}\max_{\lambda\in\mathbb{R}^{d}}(h^{T}\nu-1)\lVert\lambda\rVert_{2}+\lVert\nu\rVert_{2}g^{T}\lambda-\nu^{T}\mathbf{1},$
and when $h^{T}\nu-1\leq-\lVert g\rVert_{2}\lVert\nu\rVert_{2}$, the maximization over $\lambda$ is 0, where otherwise it is $\infty$. Hence, we can write $\phi(g,h)$ as
$\phi(g,h)=\min_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}-\nu^{T}b=-\max_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}\nu^{T}b.$
Hence, finding an almost sure upper bound of
$\max_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}\nu^{T}b$
gives an almost sure upper bound of $\Phi_{0}(X)$. Now, we know that
$\max_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}\nu^{T}b$ $=-\min_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}-\nu^{T}b$
$=-\min_{\nu\in\mathbb{R}^{d}}\max_{\begin{subarray}{c}\lambda_{1}\in\mathbb{R}_{+}^{d}\\ \mu_{1}\geq 0\end{subarray}}-\nu^{T}b-\lambda_{1}^{T}\nu+\mu_{1}(h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}-1)$
$=-\min_{\nu\in\mathbb{R}^{d}}\max_{\begin{subarray}{c}\lambda_{1}\in\mathbb{R}_{+}^{d}\\ \mu_{1}\geq 0\end{subarray}}(\mu_{1}h-\lambda_{1}-b)^{T}\nu+\mu_{1}\lVert g\rVert_{2}\lVert\nu\rVert_{2}-\mu_{1}$
$=-\max_{\begin{subarray}{c}\lambda_{1}\in\mathbb{R}_{+}^{d}\\ \mu_{1}\geq 0\end{subarray}}\min_{\nu\in\mathbb{R}^{d}}(\mu_{1}h-\lambda_{1}-b)^{T}\nu+\mu_{1}\lVert g\rVert_{2}\lVert\nu\rVert_{2}-\mu_{1}$
$=-\max_{\begin{subarray}{c}\lambda_{1}\in\mathbb{R}_{+}^{d},\mu_{1}\geq 0\\ \mu_{1}\lVert g\rVert_{2}\geq\lVert\mu_{1}h-\lambda_{1}-b\rVert_{2}\end{subarray}}-\mu_{1}.$
Hence, for any pair $\mu_{0}\geq 0,\lambda_{0}\in\mathbb{R}_{+}^{d}$ that satisfies $\mu_{0}\lVert g\rVert_{2}\geq\lVert\mu_{0}h-\lambda_{0}-b\rVert_{2}$, we know that
$\max_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{d}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}\nu^{T}b\leq\mu_{0}.$
Choose entries of $\lambda_{0}$ as
$\lambda_{0,i}=\begin{cases}c\quad if\quad h_{i}>0\\ 0\quad if\quad h_{i}<0\end{cases},$
and choose
$\mu_{0}=\frac{\lVert\lambda_{0}+b\rVert_{2}^{2}}{(\lambda_{0}+b)^{T}h}.$
Then, from Hoeffding’s inequality, we have a constant $A_{1}>0$ that satisfies with probability at least $1-\exp(-A_{1}d)$,
$0.4c^{2}d\leq\lVert\lambda_{0}\rVert_{2}^{2}\leq 0.6c^{2}d,\quad 0.3cd\leq(\sqrt{\frac{2}{\pi}}\frac{1}{2}-0.1)cd\leq\lambda_{0}^{T}h\leq(\sqrt{\frac{2}{\pi}}\frac{1}{2}+0.1)cd\leq 0.5cd.$
The two inequalities hold with high probability as they are concentration inequalities of $Ber(\frac{1}{2})$ and absolute of normal, respectively. Also, we know that there is a constant $A_{2}>0$ that satisfies with probability at least $1-\exp(-A_{2}d)$,
$-0.1cn\leq b^{T}\lambda_{0}\leq 0.1cn,\quad-0.1n\leq b^{T}h\leq 0.1n,$
since $\lVert b\rVert_{2}\leq 2\sqrt{n}$ and $h,\lambda_{0}$ are independent with $b$. At last, we have $n\leq\lVert b\rVert_{2}^{2}\leq 4n$. Combining these inequalities, with probability at least $1-\exp(-A_{1}d)-\exp(-A_{2}d)$, we have
$(0.2c+1)n\leq\lVert\lambda_{0}+b\rVert_{2}^{2}\leq(0.8c+4)n,\quad 0.2n\leq(\lambda_{0}+b)^{T}h\leq 0.6n.$
This means $\frac{c}{3}\leq\mu_{0}\leq 5c$ with probability at least $1-\exp(-A_{1}d)-\exp(-A_{2}d)$.
Also, the chosen $\mu_{0},\lambda_{0}$ are both positive. At last, we know that from the concentration inequality of chi-square distributions,
$\mathbb{P}(d-\epsilon d\leq\lVert g\rVert_{2}^{2}\leq d+\epsilon d)\geq 1-2\exp(-\frac{\epsilon^{2}}{4}d),\quad\mathbb{P}(d-\epsilon d\leq\lVert h\rVert_{2}^{2}\leq d+\epsilon d)\geq 1-2\exp(-\frac{\epsilon^{2}}{4}d).$
Hence, with probability at least $1-4\exp(-\frac{\epsilon^{2}}{4}d)$, $\lVert h\rVert_{2}^{2}-\lVert g\rVert_{2}^{2}\leq 2\epsilon d$. Now,
$\mu_{0}\lVert g\rVert_{2}\geq\lVert\mu_{0}h-\lambda_{0}-b\rVert_{2}$ $\iff(\lVert g\rVert_{2}^{2}-\lVert h\rVert_{2}^{2})\mu_{0}^{2}+2(\lambda_{0}+b)^{T}h\mu_{0}-\lVert\lambda_{0}+b\rVert_{2}^{2}\geq 0$
$\iff(\lVert g\rVert_{2}^{2}-\lVert h\rVert_{2}^{2})\mu_{0}^{2}+\lVert\lambda_{0}+b\rVert_{2}^{2}\geq 0.$
Use $\mu_{0}\leq 5c$, $\lVert\lambda_{0}+b\rVert_{2}^{2}\geq(0.2c+1)n$ to obtain that if $(\lVert h\rVert_{2}^{2}-\lVert g\rVert_{2}^{2})\leq\frac{1}{125}d$, $\mu_{0}\lVert g\rVert_{2}\geq\lVert\mu_{0}h-\lambda_{0}-b\rVert_{2}$ holds. Take $\epsilon=\frac{1}{250}$ to find $A_{3}>0$ that with probability at least $1-\exp(-A_{3}d)$, $\mu_{0},\lambda_{0}$ are pairs that satisfy $\mu_{0}\lVert g\rVert_{2}\geq\lVert\mu_{0}h-\lambda_{0}-b\rVert_{2}$, which means
$\max_{\begin{subarray}{c}\nu\in\mathbb{R}_{+}^{n}\\ h^{T}\nu+\lVert g\rVert_{2}\lVert\nu\rVert_{2}\leq 1\end{subarray}}\nu^{T}b\leq 5c,$
with probability at least $1-\exp(-A_{1}d)-\exp(-A_{2}d)-\exp(-A_{3}d)$. This result along with
$\mathbb{P}(\Phi_{0}(X)\leq-M)=\mathbb{P}(\Phi(X)\geq M)=1-\mathbb{P}(\Phi(X)\leq M)\geq 1-2\mathbb{P}(\phi(g,h)\leq M)$
when $M=-5c$ yields the proposed theorem. ∎
Theorem 11.
(Theorem 4.5. of the paper) Let $p_{c}$ be the probability that the $0.75d-th$ absolute value of $n$ samples sampled uniformly at random from $\mathcal{N}(0,1)$ is less than $d/2n=1/2c$. Also, let $p_{n}$ be the probability that $\lVert\tilde{X}z\rVert_{2}\leq 2\sqrt{n}$ and $d\geq d_{2}$, where $d_{2}$ is the threshold that we can guarantee the existence of such $\lambda\in\mathbb{R}^{d_{2}}$ from Lemma 2 with probability at least $1-\delta^{\prime\prime}$.
Then, with probability at least $1-p_{c}-p_{n}-2/(2n)^{9}-\delta^{\prime\prime}$, we know that
$C(\mathcal{K},z)\leq 1+80c^{2}\sqrt{\log(2n)},$
Here $\mathcal{K}$ is the rotated cone $\{u|\tilde{X}u\geq 0\}$, and all elements of the first column of $\tilde{X}$ are positive.
Proof.
Before proving, we can see that both $p_{c}$ and $p_{n}$ approaches zero as $n,d\rightarrow\infty$. The first property holds from the cumulative distribution of Gaussian, and the second property holds from Marchenko-Pasteur. Also, let’s rearrange the rows of $\tilde{X}$ so that the rows of $\tilde{X}$ satisfies the following:
1) For the $i$-th row $\tilde{X}_{i}$, if $|\tilde{X}_{i}\cdot z|\geq 4\sqrt{c}$, put it on the first $d/4$ rows. We know that there are at most $d/4$ such rows, as $\lVert Xz\rVert_{2}\leq 2\sqrt{n}$.
2) Then, sort the rows with the first column. Note that if $i\geq d$, $\tilde{X}[i,1]\geq 1/2c$ with high probability.
To prove the theorem, we will find appropriate $u$, $\epsilon$ for Proposition 6 to prove the theorem. First, from Lemma 2, we can choose a vector $\lambda_{0}\in\mathbb{R}^{d-1}$ that satisfies
$\lVert\lambda_{0}\rVert_{2}\leq 5c,\quad\tilde{X}[1,(d-1);2,d]\lambda_{0}\geq|\tilde{X}[1:d-1;1:d]z|,$
as $\lVert\tilde{X}[1:d-1;1:d]z\rVert_{2}\leq\lVert\tilde{X}z\rVert_{2}\leq 2\sqrt{n}$. Here, $X[a,b;c,d]$ denotes the submatrix of row $a$ to $b$, column $c$ to $d$. For that $\lambda_{0}$, take
$u=(1-\frac{1}{256c^{2}\log(2n)})\cdot e_{1}+\frac{1}{16c\sqrt{\log(2n)}}\cdot\frac{1}{5c}\cdot\begin{bmatrix}0\\ \lambda_{0}\end{bmatrix}.$
At last, take $\epsilon=\frac{1}{80c^{2}\sqrt{\log(2n)}}$. We want to see that with high probability, the chosen $u,\epsilon$ satisfies the conditions for Proposition 6. First, it is clear that $\lVert u\rVert_{2}\leq 1$, as
$\lVert u\rVert_{2}^{2}=(1-\frac{1}{256c^{2}\log(2n)})^{2}+\frac{1}{256c^{2}\log(2n)}(\frac{1}{5c}\lVert\lambda_{0}\rVert_{2})^{2}\leq(1-\frac{1}{256c^{2}\log(2n)})^{2}+\frac{1}{256c^{2}\log(2n)}\leq 1.$
The last inequality follows from $256c^{2}\log(2n)\geq 1$. By verifying the last condition, the second condition follows. Now, let’s divide the problem into cases.
If $i<d$, for the $i-th$ row $\tilde{X}_{i}$ of $\tilde{X}$,
$\tilde{X}_{i}\cdot u\geq(1-\frac{1}{256c^{2}\log(2n)})\tilde{X}[i,1]+\frac{1}{80c^{2}\sqrt{\log(2n)}}\tilde{X}[i;2,d]\cdot\lambda_{0}\geq\frac{1}{80c^{2}\sqrt{\log(2n)}}|\tilde{X}_{i}\cdot z|,$
which follows directly from the construction of $\lambda_{0}$.
If $i\geq d$, we obtain
$\tilde{X}_{i}\cdot u$ $\geq(1-\frac{1}{256c^{2}\log(2n)})\tilde{X}[i,1]+\frac{1}{80c^{2}\sqrt{\log(2n)}}\tilde{X}[i;2,d]\cdot\lambda_{0}$
$\geq(1-\frac{1}{256c^{2}\log(2n)})\frac{1}{2c}+\frac{1}{80c^{2}\sqrt{\log(2n)}}\tilde{X}[i;2,d]\cdot\lambda_{0},$
with probability $1-p_{c}$. Also, we know that the inner product between a unit vector and a random normal vector follows $\mathcal{N}(0,1)$, hence for $i\geq d$, the variables $\tilde{X}[i;2,d]\cdot\lambda_{0}$ can be understood as i.i.d. random variables sampled from $\mathcal{N}(0,\lVert\lambda_{0}\rVert_{2}^{2})$. From the concentration inequality of the maximum of absolute values, we know that
$\mathbb{P}[\max_{i\geq d}|\tilde{X}[i;2,d]\cdot\lambda_{0}|\geq 20c\sqrt{2\log(2n)}]\leq 2\cdot\frac{1}{(2n)^{9}}.$
Hence, with probability at least $1-p_{c}-\frac{2}{(2n)^{9}}$, we can see that
$\tilde{X}_{i}\cdot u$ $\geq(1-\frac{1}{256c^{2}\log(2n)})\frac{1}{2c}+\frac{1}{80c^{2}\sqrt{\log(2n)}}\tilde{X}[i;2,d]\cdot\lambda_{0}$
$\geq(1-\frac{1}{256c^{2}\log(2n)})\frac{1}{2c}-\frac{1}{4c}$
$\geq\frac{255}{512c}-\frac{1}{4c}\approx\frac{0.249}{c}\geq\frac{1}{17c\sqrt{c}}|X_{i}\cdot z|,$
for all $i\geq d$. The last inequality follows from rearrangement of the rows of $\tilde{X}$, and for $i\geq d$, $|\tilde{X}_{i}\cdot z|\leq 4\sqrt{c}$. This means that for all $i\in[n]$,
$\tilde{X}_{i}\cdot u\geq\frac{1}{80c^{2}\sqrt{\log(2n)}}|\tilde{X}_{i}\cdot z|$
holds, and $\tilde{X}u\geq\epsilon|\tilde{X}_{i}\cdot z|$. This shows that the chosen $u,\epsilon$ satisfy the last condition for Proposition 6, and the second condition. The result directly follows from Proposition 6. ∎
Corollary 7.
For subsampled hyperplane arrangement patterns $\tilde{D}_{1},\tilde{D}_{2},...,\tilde{D}_{\tilde{P}}$, let
$p_{0}^{*}=\min_{u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}),$ (17)
$p_{1}^{*}=\min_{w_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{\tilde{P}}\tilde{D}_{i}Xw_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{\tilde{P}}\lVert w_{i}\rVert_{2}.$ (18)
Furthermore, assume (A1) and $d\geq d_{2}$ as defined in Theorem 11. Then, we have $p_{1}^{*}\leq p_{0}^{*}\leq(1+80c^{2}\sqrt{\log 2n})p_{1}^{*}$ with probability at least $1-p_{c}-p_{n}-2/(2n)^{9}-\delta^{\prime\prime}$.
Proof.
The first inequality follows from triangular inequality, and the second from Proposition 5. ∎
Appendix D Proof of the Main Theorems
Theorem 12.
(Theorem 2.1. of the paper) Let the optimal value of the 2-layer ReLU network as
$p^{*}=\min_{u,\alpha}\frac{1}{2}\lVert\sum_{j=1}^{2m}(Xu_{j})_{+}\alpha_{j}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{j=1}^{2m}(\lVert u_{j}\rVert_{2}^{2}+\lVert\alpha_{j}\rVert_{2}^{2}),$
and the convex optimization problem with random hyperplane arrangement patterns as
$\tilde{p}^{*}=\min_{u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}}\frac{1}{2}\lVert\sum_{i=1}^{m/2}\tilde{D}_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{m/2}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}).$
Suppose that $m\geq\max\{m^{*},320(\sqrt{c}+1)^{2}\log(\frac{n}{\delta})\}$, where $m^{*}$ defined as in Section 2.3. Moreover, assume $n/d=c$ is fixed and the entries of $X$ are i.i.d. $\sim\mathcal{N}(0,1)$. At last, let $d\geq d_{3}$ so that both Corollary 7 and Proposition 3 holds with probability at least $1-\delta-\delta^{\prime}$ and $G<1/2$. Then,
$p^{*}\leq\tilde{p}^{*}\leq 2\sqrt{20}(\sqrt{c}+1)(1+80c^{2}\sqrt{\log(2n)})\ p^{*},$
holds with probability at least $1-\delta-\delta^{\prime}$.
Proof.
Let’s denote
$p_{2}^{*}=\min_{u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}}\frac{1}{2}\lVert\sum_{i=1}^{P}D_{i}X(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{P}(\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}),$
$p_{3}^{*}=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{P}D_{i}Xu_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{P}\lVert u_{i}\rVert_{2},$
where $D_{1},D_{2},...,D_{P}$ are all possible hyperplane arrangement patterns, and
$p_{4}^{*}=\min_{u_{i}\in\mathbb{R}^{d}}\frac{1}{2}\lVert\sum_{i=1}^{m/2}\tilde{D}_{i}Xu_{i}-y\rVert_{2}^{2}+\beta\sum_{i=1}^{m/2}\lVert u_{i}\rVert_{2},$
where $\tilde{D}_{1},\tilde{D}_{2},...,\tilde{D}_{m/2}$ are randomly sampled hyperplane arrangement patterns. First, from preliminaries, we know that $p^{*}=p_{2}^{*}$. It is clear that $p_{2}^{*}\leq\tilde{p}^{*}$, as we use less hyperplane arrangement patterns during approximation. Moreover, we know that $p_{3}^{*}\leq p_{2}^{*}$, as $\lVert u_{i}-v_{i}\rVert_{2}\leq\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}$ and $u_{i}-v_{i}$ can represent arbitrary vector in $\mathbb{R}^{d}$ even with the constraint $u_{i},v_{i}\in\mathcal{K}_{\tilde{D}_{i}}$. To wrap up, we know that
$p_{3}^{*}\leq p_{2}^{*}=p^{*}\leq\tilde{p}^{*}.$
From Corollary 7 we know that $\tilde{p}^{*}\leq(1+80c^{2}\sqrt{\log 2n})\ p_{4}^{*}$, and we also know from Proposition 4 that $p_{3}^{*}\geq G\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}$. At last, from Proposition 3, we know that $p_{4}^{*}\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}}$. Hence, we know that
$G\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{\max}(XX^{T})}}\leq p_{3}^{*}\leq p_{2}^{*}=p^{*}\leq\tilde{p}^{*}$ $\leq(1+80c^{2}\sqrt{\log 2n})\ p_{4}^{*}$
$\leq\sqrt{2}\beta\frac{\lVert y\rVert_{2}}{\sqrt{\lambda_{min}(\mathcal{M})}}\cdot(1+80c^{2}\sqrt{\log 2n})$
$\leq\frac{\sqrt{2}}{G}\sqrt{\frac{\lambda_{\max}(XX^{T})}{\lambda_{\min}(\mathcal{M})}}\cdot(1+80c^{2}\sqrt{\log 2n})\ p^{*}$
$\leq\frac{\sqrt{20}}{G}(\sqrt{c}+1)(1+80c^{2}\sqrt{\log 2n})\ p^{*},$
which finishes the proof that
$p^{*}\leq\tilde{p}^{*}\leq 2\sqrt{20}(\sqrt{c}+1)(1+80c^{2}\sqrt{\log(2n)})\ p^{*}.$
Theorem 13.
[…]
Proof.
Consider the Gaussian relaxation of the convex reformulation with $[m/2]$ hyperplane arrangement patterns. As there are $O(dm)$ variables, we can solve the problem with $O(d^{3}m^{3})$ complexity using standard interior point solvers. Moreover, we have the approximation bound of Theorem 1, which leads to the fact that the solved global minima has $O(\sqrt{\log(2n)})$ guarantees. At last, we can map the solution $\{(u_{i}^{*},v_{i}^{*})\}_{i=1}^{[m/2]}$ to the parameter space of two-layer neural networks with the mapping
$u_{i}^{*}\rightarrow(\frac{u_{i}^{*}}{\sqrt{\lVert u_{i}^{*}\rVert_{2}}},\sqrt{\lVert u_{i}^{*}\rVert_{2}}),v_{i}^{*}\rightarrow(\frac{v_{i}^{*}}{\sqrt{\lVert v_{i}^{*}\rVert_{2}}},-\sqrt{\lVert v_{i}^{*}\rVert_{2}}),$
to find the parameters of the two-layer neural network that has the same loss function value as the optimal value of the convex problem. Hence, we can find parameters of two-layer neural network that has $O(\sqrt{\log(n)})$ relative optimality bound in $O(d^{3}m^{3})$ time. ∎
Theorem 14.
(Theorem 2.5. of the paper) Consider the training problem $\min_{u_{j},\alpha_{j}}\mathcal{L}(u,\alpha),$ where the loss function $\mathcal{L}$ is given as
$\mathcal{L}(u,\alpha)=\frac{1}{2}\lVert\sum_{j=1}^{m}(Xu_{j})_{+}\alpha_{j}-y\rVert_{2}^{2}+\frac{\beta}{2}\sum_{j=1}^{m}(\lVert u_{j}\rVert_{2}^{2}+\alpha_{j}^{2}).$
Assume (A1),(A2) and $d\geq d_{3}$ so that Corollary 7 holds with probability $1-\delta$, and $m\geq\max{m^{*},320(\sqrt{c}+1)^{2}\log(\frac{n}{\delta^{\prime}})}$. For any random initialization $\{u_{i}^{0},\alpha_{i}^{0}\}_{i=1}^{m}$, suppose local gradient method converged to a stationary point $\{u_{i}^{\prime},\alpha_{i}^{\prime}\}_{i=1}^{m}$. Then, with probability at least $1-\delta-\delta^{\prime}$,
$\mathcal{L}(u^{\prime},\alpha^{\prime})\leq C\sqrt{\log(2n)}\mathcal{L}(u^{*},\alpha^{*}),$
for some $C\geq 1$. Here, $\{u_{i}^{*},\alpha_{i}^{*}\}_{i=1}^{m}$ is a global optimum of $\mathcal{L}(u,\alpha)$.
Proof.
We know that for the stationary point $\{u_{i}^{\prime},\alpha_{i}^{\prime}\}_{i=1}^{m}$, we have a corresponding convex optimization problem
$p^{*}:=\frac{1}{2}\lVert\sum_{i=1}^{m}\tilde{D}_{i}(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{m}\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2},$ (19)
where $\tilde{D}_{i}=\mathbbm{1}(Xu_{i}^{\prime}\geq 0)$ and the solution mapping of the optimal solution $\{(u_{i}^{*},v_{i}^{*})\}_{i=1}^{m}$ of the convex problem and the stationary point given as
$\alpha_{i}^{\prime}=sign(\alpha_{i}^{\prime})\sqrt{\lVert u_{i}^{*}\rVert_{2}},\quad u_{i}^{\prime}=\frac{u_{i}^{*}}{\sqrt{\lVert u_{i}^{*}\rVert_{2}}}\mathbbm{1}(\alpha_{i}^{\prime}\geq 0)+\frac{v_{i}^{*}}{\sqrt{\lVert v_{i}^{*}\rVert_{2}}}\mathbbm{1}(\alpha_{i}^{\prime}<0).$
By (A2), we know that at least half of the hyperplane arrangement patterns are preserved from $\mathbbm{1}(Xu_{0}\geq 0)$. Let the hyperplane arrangement patterns be $\tilde{D}_{1},\tilde{D}_{2},...,\tilde{D}_{m/2}$ without loss of generality. Now, we know that the optimal solution of the problem
$p^{\prime}:=\frac{1}{2}\lVert\sum_{i=1}^{m/2}\tilde{D}_{i}(u_{i}-v_{i})-y\rVert_{2}^{2}+\beta\sum_{i=1}^{m/2}\lVert u_{i}\rVert_{2}+\lVert v_{i}\rVert_{2}$
has $O(\sqrt{\log 2n})$ approximation guarantee with probability at least $1-\delta-\delta^{\prime}$, and as it is using less hyperplane arrangement patterns, we know that $p^{*}\leq p^{\prime}$. At last, we know that $p^{*}=\mathcal{L}(u^{\prime},\alpha^{\prime})$ from solution mapping. Plugging in $\mathcal{L}(u^{\prime},\alpha^{\prime})$ at $p^{*}$ and using Theorem 1 on $p^{\prime}$ yields the wanted result. ∎